### v2

# Any-Step Density Ratio Estimation via Interval‑Annealed Secant Alignment

Abstract
Estimating density ratios is a fundamental problem in machine learning, but existing methods often trade off accuracy for efficiency. We propose Interval-annealed Secant Alignment Density Ratio Estimation (ISA-DRE), a framework that enables accurate, any-step estimation without numerical integration. Instead of modeling infinitesimal tangents as in prior methods, ISA-DRE learns a global secant function, defined as the expectation of all tangents over an interval, with provably lower variance, making it more suitable for neural approximation. This is made possible by the Secant Alignment Identity, a self-consistency condition that formally connects the secant with its underlying tangent representations. To mitigate instability during early training, we introduce Contraction Interval Annealing, a curriculum strategy that gradually expands the alignment interval during training. This process induces a contraction mapping, which improves convergence and training stability. Empirically, ISA-DRE achieves competitive accuracy with significantly fewer function evaluations compared to prior methods, resulting in much faster inference and making it well suited for real-time and interactive applications.
1 Introduction
Estimating the density ratio, $r(\mathbf{x})=p_{1}(\mathbf{x})/p_{0}(\mathbf{x})$, is a core problem in machine learning, supporting key applications in domain adaptation (Wang et al. 2023), causal inference (Wang et al. 2025) and $f$-divergence estimation (Chen et al. 2025). A common approach is to estimate $p_{0}$ and $p_{1}$ separately, but this becomes inefficient or unstable when the densities are intractable or significantly different–a challenge known as the density-chasm problem.
Existing methods like noise-contrastive estimation (Gutmann and Hyvärinen 2010) and trimmed estimators (Liu et al. 2017) offer partial relief but suffer from high variance or intensive hyperparameter tuning. To bridge this gap, TRE (Rhodes, Xu, and Gutmann 2020) decomposes the ratio into a product over interpolated intermediate distributions. While effective against density chasms, TRE requires training $M$ separate models and remains sensitive to large distribution shifts (Choi et al. 2022; Chen et al. 2025).
In the limit $M\to\infty$, the tangent-based method, DRE-$\infty$ (Choi et al. 2022), reformulates the ratio as a time integral:
$\log r(\mathbf{x})=\int_{0}^{1}\partial_{\tau}\log p_{\tau}(\mathbf{x})\mathrm{d}{\tau},$
enabling single-model training via score matching on the time score function $\partial_{\tau}\log p_{\tau}$. While this avoids ensembling, it incurs heavy inference costs from numerical quadrature–often requiring hundreds of evaluations per sample.
Subsequent methods improve stability via diffusion-bridge interpolants (Chen et al. 2025) and improve efficiency via conditional time score matching (Yu et al. 2025), but the reliance on numerical integration remains a key bottleneck.
In this work, we propose a fundamentally different method. Rather than estimating the infinitesimal time score $\partial_{\tau}\log p_{\tau}$ (i.e., the tangent function), we directly learn its integral, which is termed as the secant function:
$u(\mathbf{x},l,t)=\frac{1}{t-l}\int_{l}^{t}\partial_{\tau}\log p_{\tau}(\mathbf{x})\mathrm{d}\tau,\ l<t,$
which captures the average change in log-density over the interval $l,t]\subseteq[0,1]$. Notably, the desired log-density ratio naturally arises as a special case:
$\log r(\mathbf{x})=u(\mathbf{x},0,1)=\int_{0}^{1}\partial_{\tau}\log p_{\tau}(\mathbf{x})\mathrm{d}{\tau}.$
By learning the secant directly, we bypass numerical integration entirely, enabling stable and efficient density ratio estimation (DRE) even under large discrepancies.
We propose Interval‑annealed Secant Alignment Density Ratio Estimation (ISA-DRE), a framework that reformulates DRE as a direct function approximation problem over secant intervals, thereby eliminating the need for numerical solvers or quadrature. At its core is the Secant Alignment Identity, which provides a principled bridge between secant and tangent representations. We theoretically show that minimizing this loss retains the consistency guarantees of tangent-based methods, while benefiting from reduced variance and improved training stability. By replacing costly integration with any-step inference, ISA-DRE offers a practical solution to the long-standing trade-off between accuracy and efficiency. Empirically, it matches the accuracy of state-of-the-art methods with far fewer function evaluations, making it particularly effective for real-time and interactive applications.
2 Related Works
Classical Density Ratio Estimation (DRE).
Seminal works in DRE follow two main avenues: estimating each density separately (e.g., via kernel density estimation (Huang et al. 2006)) or modeling the ratio directly by minimizing a statistical divergence, as in KLIEP (Sugiyama et al. 2008) and uLSIF (Kanamori, Hido, and Sugiyama 2009). Discriminative approaches such as NCE (Gutmann and Hyvärinen 2012) and TR-DRE (Liu et al. 2017) have been widely adopted, with further theoretical insights into truncation given by Song and Ermon (2019). Despite their simplicity, these methods often fail under the density-chasm problem, where $p_{0}$ and $p_{1}$ have little or no overlap, leading to unstable training and poor generalization in high-dimensional settings. To bridge this gap, divide-and-conquer strategies have been proposed. TRE (Rhodes, Xu, and Gutmann 2020) decomposes the ratio into a sequence of intermediate steps, each estimated separately, and AM-DRE (Xiao and Han 2022) adaptively refines this idea for latent-space EBM learning. Their continuous counterpart, DRE-$\infty$ (Choi et al. 2022), introduces a time-dependent score function defined along a continuous interpolation path, marking a conceptual shift from static ratios to dynamical estimation. However, these methods rely on expensive numerical integration, often requiring hundreds of evaluations per sample. Subsequent refinements explored alternative parameterizations and regularization schemes: FDRE (Choi, Liao, and Ermon 2021) learns featurized ratios to improve generalization; IMDRE (Kimura and Bondell 2025) combines importance sampling with multi-step estimation; D3RE (Chen et al. 2025), $\gamma$-DRE (Nagumo and Fujisawa 2024), and iterated regularization (Gruber et al. 2024) further enhance robustness under large-support mismatch. Srivastava et al. (2023) reformulate DRE as multinomial logistic regression to better handle high-discrepancy distributions, and Williams et al. (2025) exploit exponential-family representations to reduce variance. Despite these advances, the computational bottleneck posed by numerical integration persists, motivating approaches like ours that preserve the continuity and robustness of DRE-$\infty$ while eliminating solver overhead.
Any-Step Inference.
In the parallel domain of generative modeling, consistency models (Song et al. 2023) and progressive distillation (Salimans and Ho 2022) have enabled efficient single-step generation by training a student network to replicate the outcome of multi-step inference, effectively collapsing the inference pipeline. However, their reliance on pre-trained teacher models or complex multi-stage curricula renders them ill-suited for density ratio estimation, where ground-truth ratios for distillation are unavailable.
Our Contribution: ISA-DRE.
ISA-DRE eliminates the need for both numerical solvers and teacher-student distillation. Instead of learning instantaneous changes (tangents) and integrating them afterward, it directly learns average changes over intervals (secants). By enforcing the Secant Alignment Identity through interval annealing, ISA-DRE enables direct estimation of the density ratio in an any-step fashion, without relying on numerical solvers or auxiliary teacher models.
3 Background
Let $p_{0}(\mathbf{x})$ and $p_{1}(\mathbf{x})$ be two probability density functions. The objective of DRE is to estimate the ratio $r(\mathbf{x})=p_{1}(\mathbf{x})/p_{0}(\mathbf{x})$ from samples drawn from these distributions, without access to their analytical forms. A principal challenge arises when the supports of $p_{0}$ and $p_{1}$ are largely disjoint (Gutmann and Hyvärinen 2012; Liu et al. 2017).
To address this, path-based methods construct a family of intermediate distributions $\{p_{t}\}_{t\in[0,1]}$ that smoothly connect $p_{0}$ and $p_{1}$. TRE (Rhodes, Xu, and Gutmann 2020) introduced this idea with a discrete “divide-and-conquer” strategy, using $M$ intermediate steps based on a linear interpolant,
$\mathbf{X}_{m/M}=\sqrt{1-\beta_{m/M}^{2}}\mathbf{X}_{0}+\beta_{m/M}\mathbf{X}_{1},$ (1)
where $\mathbf{X}_{0}\sim p_{0}$, $\mathbf{X}_{1}\sim p_{1}$, and $\{\beta_{m/M}\}$ is an increasing sequence. The total ratio is then a telescoping product of intermediate ratios, $r(\mathbf{x})=\prod_{m=0}^{M-1}r_{m/M}(\mathbf{x}).$ While theoretically sound, this method necessitates training $M$ separate networks, posing a significant computational burden and leaving the density chasm partially unmitigated with small $M$.
DRE-$\infty$ (Choi et al. 2022) generalized this approach to the continuous limit ($M\to\infty$), defining a continuous path via a deterministic interpolant (DI),
$\mathbf{X}_{t}=\alpha_{t}\mathbf{X}_{0}+\beta_{t}\mathbf{X}_{1}\quad\text{for }t\in[0,1],$ (2)
where coefficients $\alpha_{t},\beta_{t}$ ensure that $\mathbf{X}_{t}$ smoothly transitions from being distributed as $p_{0}$ at $t=0$ to $p_{1}$ at $t=1$. The log-density ratio is then estimated via:
$\log r(\mathbf{x})=\int_{0}^{1}\partial_{\tau}\log p_{\tau}(\mathbf{x})\mathrm{d}{\tau}=\int_{0}^{1}s_{t}(\mathbf{x},\tau)\mathrm{d}{\tau},$ (3)
where $s_{t}(\mathbf{x},\tau)\triangleq\partial_{\tau}\log p_{\tau}(\mathbf{x})$ is the time score function. To enhance robustness and stability, D3RE (Chen et al. 2025) proposed the dequantified diffusion-bridge interpolant (DDBI), which stabilizes the path with Gaussian noise,
$\mathbf{X}_{t}=\alpha_{t}\mathbf{X}_{0}+\beta_{t}\mathbf{X}_{1}+\sqrt{t(1-t)\gamma^{2}+(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon}\mathbf{Z},$ (4)
where $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, $\gamma\in\mathbb{R}_{\geq 0}$ is the noise factor, and $\varepsilon$ is a small number for stability.
In these frameworks, a score model $s^{\boldsymbol{\theta}}_{t}$ is trained to approximate the true score $s_{t}$ by minimizing a time score-matching (TSM) loss (Choi et al. 2022). Since the marginal score $s_{t}$ is intractable, training instead uses the equivalent conditional time score-matching (CTSM) objective (Yu et al. 2025),
$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}\left[\lambda(t)\left|s_{t}(\mathbf{x}_{t},t\mid\mathbf{y})-s^{\boldsymbol{\theta}}_{t}(\mathbf{x}_{t},t)\right|^{2}\right],$ (5)
where $\lambda(t)\propto 1/\text{Var}_{p_{t}}(s_{t}\mid\mathbf{y})$ is a weighting function, the conditioning variable $\mathbf{y}$ is $\mathbf{x}_{1}$ for DI and $(\mathbf{x}_{0},\mathbf{x}_{1})$ for DDBI, and the conditional score $s_{t}(\mathbf{x}_{t},t\mid\mathbf{y})=\partial_{t}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{y})$ is tractable. After training, the ratio is estimated by numerically integrating the learned score: $\log r^{\boldsymbol{\theta}^{\star}}(\mathbf{x})=\int_{0}^{1}s^{\boldsymbol{\theta}^{\star}}_{t}(\mathbf{x},t)\mathrm{d}t$.
Figure 1: A geometric illustration of tangent- vs. secant-based density ratio estimation. Conventional methods estimate $\log r(\mathbf{x})$ by numerically integrating the tangent function (blue curve). In contrast, ISA-DRE learns a secant function, whose value at $\tau=\xi$ equals the average height over the interval, making the red rectangle’s area exactly match the area under the blue curve. This provides a more direct and efficient approach to estimating $\log r(\mathbf{x})$.
4 Methods
This section details the underlying theory and practical implementation of our ISA-DRE framework.
4.1 Secant Alignment
The Secant Function.
Our approach is based on the secant function $u$, which is the average of the time score function $s_{t}$–also called the tangent function–over a time interval $[l,t]$,
[…]
In the limiting case $t=l$, we define $u(\mathbf{x}_{t},t,t)=s_{t}(\mathbf{x}_{t},t)$ since $u$ is continuous at $t=l$, i.e.,
$u(\mathbf{x}_{t},t,t)=\lim_{l\to t}\frac{1}{t-l}\int_{l}^{t}s_{t}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau=s_{t}(\mathbf{x}_{t},t).$ (7)
This definition reveals the relationship between the secant and the tangent functions:
$u(\mathbf{x}_{t},l,t)=\mathbb{E}_{p(\tau)}\left[s_{t}(\mathbf{x}_{\tau},\tau)\right]=(t-l)\cdot s_{t}(\mathbf{x}_{\xi},\xi),$ (8)
where $p(\tau)=\mathcal{U}[l,t]$ and $\xi\in[l,t]$. The first equality follows by definition, and the second by the mean value theorem for integrals, assuming $s_{t}$ is continuous on $l,t]\subseteq[0,1]$.
The equations in Eq. 8 reveal two key insights:
  • (1)
    The secant over the interval $[l,t]$ represents the expectation of all tangents between $l$ and $t$;
  • (2)
    The secant can be interpreted as a reparameterized tangent, evaluated at some intermediate point $\xi\in[l,t]$.
A geometric illustration of (2) is shown in Fig. 1. Building on (1), the secant exhibits lower variance than the tangent (see Proposition 4.1), making it more suitable for approximation.
Proposition 4.1.
Let $l$ and $t$ be independent random variables with probability density $p(\cdot)$ on $[0,1]$, conditioned on $l\leq t$. For a fixed data point $\mathbf{x}\sim p_{1}$, define the secant variable $U\triangleq u(\mathbf{x},l,t)$ and tangent variable $S\triangleq s_{t}(\mathbf{x},\tau)$, where $\tau\sim p(\tau)=\mathcal{U}[l,t]$. Under the joint distribution $p(l,t)$, the variance of $U$ w.r.t. $(l,t)$ satisfies:
$\text{Var}_{p(l,t)}(U)\leq\text{Var}_{p(\tau)}(S),$ (9)
with equality $\mathrm{iff}$ $S$ is constant for $p$-almost every $\tau\in[0,1]$.
The Secant Alignment Identity (SAI).
To enable learning of the secant function without explicitly computing time integrals, we derive a differential relationship that links the secant to the tractable conditional tangent function.
Unlike the simplified setting used in DRE, where $\mathbf{x}_{t}\equiv\mathbf{x}$ for all $t\in[0,1]$, we now consider a general time-dependent interpolant $\mathbf{x}_{t}\sim p_{t}$ that varies with $t$. By rearranging the integral definition of $u$ as $(t-l)u(\mathbf{x}_{t},l,t)=\int_{l}^{t}s_{t}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau$ and differentiating both sides w.r.t. $t$, we obtain the Secant Alignment Identity (SAI),
$\underbrace{u(\mathbf{x}_{t},l,t)}_{\text{Secant Function}}=\underbrace{s_{t}(\mathbf{x}_{t},t)}_{\text{Tangent Function}}-\underbrace{(t-l)\cdot\frac{\mathrm{d}}{\mathrm{d}t}u(\mathbf{x}_{t},l,t)}_{\text{Correction Term}},$ (10)
where $\frac{\mathrm{d}}{\mathrm{d}t}u$ is the total derivative of $u$ w.r.t. $t$. Since $\frac{\mathrm{d}l}{\mathrm{d}t}=0$ and $\frac{\mathrm{d}t}{\mathrm{d}t}=1$, this derivative can be derived via the chain rule:
$\frac{\mathrm{d}}{\mathrm{d}t}u(\mathbf{x}_{t},l,t)$ $=\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}\cdot\partial_{\mathbf{x}}u+\frac{\mathrm{d}l}{\mathrm{d}t}\cdot\partial_{l}u+\frac{\mathrm{d}t}{\mathrm{d}t}\cdot\partial_{t}u$ (11)
$=\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}\cdot\partial_{\mathbf{x}}u+\partial_{t}u.$
Here, the derivative term $\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$ is known analytically from the chosen interpolant, e.g., Eq. 2 or Eq. 4,
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\begin{cases}\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}\mathbf{x}_{0}+\frac{\mathrm{d}\beta_{t}}{\mathrm{d}t}\mathbf{x}_{1},&\text{ if DI},\\ \frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}\mathbf{x}_{0}+\frac{\mathrm{d}\beta_{t}}{\mathrm{d}t}\mathbf{x}_{1}+\frac{\gamma(1-2t)}{2\sqrt{t(1-t)}}\mathbf{z},&\text{ if DDBI},\end{cases}$ (12)
where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, and $\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$, $\frac{\mathrm{d}\beta_{t}}{\mathrm{d}t}$ denote time derivatives of the coefficients. The partial derivatives $\partial_{\mathbf{x}}u$ and $\partial_{t}u$ are evaluated using the Jacobian-vector product (JVP) between the Jacobian $\partial_{\mathbf{x}}u,\partial_{l}u,\partial_{t}u]$ and the direction vector $\left[\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t},0,1\right]$.
The SAI thus provides a self-consistency condition: the ground-truth secant function over the interval $[l,t]$ can be reconstructed from the typically modeled tangent at $t$ minus a correction term. This relationship forms the learning framework for training the tangent $u$.
Training with the SAI.
The Proposition 4.1 and the SAI provide a clear prescription for training a neural network $u^{\boldsymbol{\theta}}$ to approximate the true secant function $u$.
We replace the parameterized tangent term in the CTSM objective (Eq. 5) with its SAI-based representation. This yields our Conditional Secant Alignment (CSA) loss:
$\mathcal{L}_{\text{CSA}}(\boldsymbol{\theta})=\mathbb{E}\left[\lambda(t)\left|s_{t}(\mathbf{x}_{t},t\mid\mathbf{y})-s^{\boldsymbol{\theta}}_{t}(\mathbf{x}_{t},t)\right|^{2}\right],$ (13)
where $s^{\boldsymbol{\theta}}_{t}(\mathbf{x}_{t},t)=u^{\boldsymbol{\theta}}(\mathbf{x}_{t},l,t)+\text{sg}\left((t-l)\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}}(\mathbf{x}_{t},l,t)\right)$ with sg being the stop-gradient operation following common practice (Song and Dhariwal 2024). This expectation is taken over time pair $(l,t)$, sample pair $(\mathbf{x}_{0},\mathbf{x}_{1})$, the interpolant $\mathbf{x}_{t}$ and conditioning variable $\mathbf{y}$. By minimizing this loss, the network $u^{\boldsymbol{\theta}}$ is trained to be consistent with the underlying secants across all possible sub-intervals. More details for training and inference are presented in Algorithm 1.
Algorithm 1 Training and inference procedures for ISA-DRE
Input: Secant model $u^{\boldsymbol{\theta}}$, data distributions $p_{0},p_{1}$, time distribution $p(l,t)$, interpolant schedules.
Output: trained secant model $u^{\boldsymbol{\theta}^{\star}}$.
1: // Training procedure for ISA-DRE
2:repeat
3:  Sample a time pair using CIA: $l,t\sim p(l,t)$.
4:  Sample a sample pair $(\mathbf{x}_{0},\mathbf{x}_{1})\sim p_{0}\times p_{1}$.
5:  Construct the interpolant point $\mathbf{x}_{t}$ using $(\mathbf{x}_{0},\mathbf{x}_{1})$.
6:  Set conditional variable $\mathbf{y}$ via interpolant schedules.
7:  Evaluate target conditional tangent $s_{t}\leftarrow s_{t}(\mathbf{x}_{t},t\mid\mathbf{y})$.
8:  Estimate secant $u^{\boldsymbol{\theta}}\leftarrow u^{\boldsymbol{\theta}}(\mathbf{x}_{t},l,t)$.
9:  Estimate $\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}}\leftarrow\text{JVP}(u^{\boldsymbol{\theta}},(\mathbf{x}_{t},l,t),(\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t},0,1))$.
10:  Estimate tangent $s_{t}^{\boldsymbol{\theta}}\leftarrow u^{\boldsymbol{\theta}}+\text{sg}\left((t-l)\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}}\right)$.
11:  Compute loss $\mathcal{L}_{\text{CSA}}\leftarrow|s_{t}-s_{t}^{\boldsymbol{\theta}}|^{2}$.
12:  Update parameters $\boldsymbol{\theta}$ using $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{CSA}}$.
13:until convergence
14: // Inference procedure for ISA-DRE
15: Sample a data $\mathbf{x}\sim p_{1}$.
16: Estimate log density ratio $\log r^{\boldsymbol{\theta}^{\star}}(\mathbf{x})=u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},0,1)$.
The validity of minimizing the CSA loss is supported by the following theoretical guarantee, which ensures that our objective uniquely recovers the true secant:
Proposition 4.2 (Secant-Tangent Consistency Guarantee).
Let the time score function, $s_{t}$, be continuous in $t$, and the secant model $u^{\boldsymbol{\theta}}$ be continuously differentiable in $t$. Then, the learned secant $u^{\boldsymbol{\theta}^{\star}}$ exactly matches the true secant:
$u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)=u(\mathbf{x},l,t)=\frac{1}{t-l}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau,$ (14)
for all $(\mathbf{x},l,t)$ with $l\neq t$, if and only if the following hold: (1) Boundary condition: $\lim_{t\to t_{0}}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},t_{0},t)=s_{t}(\mathbf{x},t_{0})$ for any fixed $t_{0}\in[0,1]$. (2) Consistency condition (SAI): $s_{t}(\mathbf{x},t)=u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)+(t-l)\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)$.
This result shows that enforcing the SAI is not just a training heuristic but a necessary and sufficient condition for recovering the exact secant. By satisfying conditions (1) and (2), ISA-DRE ensures that $u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},0,1)$ produces the correct $\log r(\mathbf{x})$ at inference time.
Any‐Step Density Ratio Estimation.
By the Fundamental Theorem of Calculus, the secant function $u$ satisfies $\log p_{t}(\mathbf{x})-\log p_{l}(\mathbf{x})=(t-l)u(\mathbf{x},l,t)$. In particular, for a given $\mathbf{x}\sim p_{1}$, choosing $l=0$ and $t=1$ gives a one‐step estimator for the desired log-density ratio,
$\log r(\mathbf{x})=\log p_{1}(\mathbf{x})-\log p_{0}(\mathbf{x})=u(\mathbf{x},0,1).$ (15)
Once $u^{\boldsymbol{\theta}^{\star}}$ is learned, the log-ratio can be estimated by:
$\log r^{\boldsymbol{\theta}^{\star}}(\mathbf{x})=u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},0,1),\mathbf{x}\sim p_{1}.$ (16)
More generally, the log‐ratio can be estimated in $K\in\mathbb{Z}_{+}$ steps by partitioning $[0,1]$ into $0=t_{0}<t_{1}<\cdots<t_{K}=1$ and applying the finite additivity of the Riemann integral:
$\log r(\mathbf{x})$ $=\sum_{k=0}^{K-1}\left(\log p_{t_{k+1}}(\mathbf{x})-\log p_{t_{k}}(\mathbf{x})\right)$ (17)
$=\sum_{k=0}^{K-1}(t_{k+1}-t_{k})u(\mathbf{x},t_{k},t_{k+1}).$
Here $K$ corresponds to the number of function evaluations (NFE) as defined in Chen et al. (2018). This property lets ISA-DRE bypass costly ODE solvers or quadrature schemes: after training, density‐ratio inference requires exactly NFE network evaluations and avoids recursive integration errors, enabling efficient, real‐time DRE.
4.2 Contraction Interval Annealing
Bootstrap Divergence.
The CSA objective uses a secant approximation $u^{\theta}$ whose target involves its own time derivative. When the interval length |tl||t-l| is large, the term $(t-l)\frac{\mathrm{d}}{\mathrm{d}t}u^{\theta}$ amplifies the initially noisy derivative $\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}}$, causing bootstrap divergence in early training. This divergence is empirically evident in Fig. 2, which compares mutual information estimation under different training settings.
Figure 2: Mutual information estimation on the Additive Noise dataset with CIA and fixed tangent ratios. The tangent ratio denotes the proportion of samples with $l=t$, corresponding to tangent-only ($100\%$) or secant-only ($0\%$) supervision (see Sec. 4.3). Shaded areas show “std” across samples. CIA ensures stable and consistent convergence.
Formally, defining the learning operator via SAI:
$\mathcal{T}(u)\triangleq s_{t}-(t-l)\frac{\mathrm{d}}{\mathrm{d}t}u.$ (18)
Training aims to find a fixed point $u$ of this operator, i.e., a function satisfying $u=\mathcal{T}(u)$. By the Banach fixed-point theorem, fixed-point iteration converges only if
$\|\mathcal{T}(u_{1})-\mathcal{T}(u_{2})\|\leq C\|u_{1}-u_{2}\|,\text{for some }C<1.$ (19)
In a normed space, where the norm of the derivative operator $\left\|\frac{\mathrm{d}}{\mathrm{d}t}\right\|$ is well-defined, we have
$\|\mathcal{T}(u_{1})-\mathcal{T}(u_{2})\|$ $=|t-l|\left\|\frac{\mathrm{d}}{\mathrm{d}t}(u_{1}-u_{2})\right\|$ (20)
$\leq|t-l|\cdot\left\|\frac{\mathrm{d}}{\mathrm{d}t}\right\|\cdot\|u_{1}-u_{2}\|,$
so the contraction constant is $C=|t-l|\cdot\left\|\tfrac{\mathrm{d}}{\mathrm{d}t}\right\|$. However, in common spaces such as $L^{2}([0,1])$, the derivative operator is unbounded, so large |tl||t-l| typically violate $C<1$, explaining the observed bootstrap divergence and instability.
Contraction Interval Annealing (CIA).
We propose CIA, a curriculum that bounds the secant interval length $d_{\max}=|t-l|$ early in training and then anneals it to 1. By starting with $d_{\max}\to 0$, the correction term
$\left\|(t-l)\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}}\right\|\leq d_{\text{max}}\left\|\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}}\right\|\to 0,$ (21)
reduces CSA to local tangent alignment: $u^{\boldsymbol{\theta}}(\mathbf{x}_{t},l,t)\approx s_{t}^{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$, and guarantees the contraction constant $C\to 0$. As training stabilizes, the derivative $\left\|\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}}\right\|$ is reliably estimated and $d_{\max}$ is gradually increased to $1$, enabling secant supervision while avoiding $C$ being large during training.
4.3 Practical Choices
We detail the practical setup of ISA-DRE. Implementation details and ablation studies are given in Sec. 5.4.
Time Sampler for Interval Sampling.
Secant intervals $(l,t)$ are drawn by three schemes: (1) Uniform (Uni.), where $l,t\sim\mathcal{U}(0,1)$; (2) Logit‑Normal (LN) via a logistic map on Gaussian samples (Esser et al. 2024; Geng et al. 2025); (3) Variance‑based Importance (VI), with $t\sim p(t)\propto 1/\mathrm{Var}_{p_{t}}(s_{t}\mid\mathbf{y})$ (see Eq. 5). Each pair $(l,t)$ is then ordered so that $l\leq t$, ensuring valid forward intervals.
Secant-Tangent Supervision (STS).
We adopt a fixed supervision ratio following Geng et al. (2025), where a fixed proportion of training pairs use the tangent case ($l=t$) and the rest use secant ($l\neq t$). The two extremes, $0\%$ and $100\%$, correspond to secant-only and tangent-only supervision, respectively. In contrast, CIA employs an adaptive curriculum by gradually increasing the maximum interval length |tl||t-l|.
5 Experiments
[…]
5.1 Illustration of the Secant vs. the Tangent
To compare the learned secant and tangent functions, we visualize their trajectories over time in Fig. 3. The left panel shows the secant $u(\mathbf{x},0,t)$ and the right shows the tangent $u(\mathbf{x},t,t)$, with each orange curve representing one fixed $\mathbf{x}$.
The secant curves are smoother and more concentrated, especially at early timesteps, while the tangent curves fluctuate more and show greater variance. Besides, secants cluster around their mean, whereas tangents are more spread out. This highlights the stability and learnability advantages of ISA-DRE, as detailed in Proposition 4.1.
Figure 3: Comparison of the learned secant function $u(\mathbf{x},0,t)$ (left) and tangent function $u(\mathbf{x},t,t)$ (right). Each orange curve shows $u$ over time $t$ for a fixed $\mathbf{x}$. The secant curves are smoother and more concentrated.
5.2 Density Estimation
To assess our model’s ability to capture the true data distribution $p_{\mathrm{data}}$, we perform density estimation by expressing complex, possibly intractable data distribution $p_{\text{data}}$ in terms of a simple base distribution $p_{0}$ (e.g., $\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$) using the density ratio $r(\mathbf{x})=p_{\text{data}}(\mathbf{x})/p_{0}(\mathbf{x})$. This yields $\log p_{\mathrm{data}}(\mathbf{x})=\log r(\mathbf{x})+\log p_{0}(\mathbf{x})$, so that density estimation reduces to approximating $r(\mathbf{x})$, and the log‐likelihood is estimated via $\log p_{\mathrm{data}}(\mathbf{x})\approx\log r^{\boldsymbol{\theta}^{\star}}(\mathbf{x})+\log p_{0}(\mathbf{x})$. All three methods employ DI (see Eq. 2), with linear schedules for tabular datasets and VP schedules (Song et al. 2021) for others, following the setup in Chen et al. (2025).
Structured and Multi‑modal Datasets.
[…]
NFE Method TS STS POWER GAS HEPMASS MINIBOONE BSDS300
10 DRE-$\infty$ VI $100\%$ $0.03\pm 0.17$ $4.34\pm 0.60$ $20.43\pm 0.52$ $20.57\pm 0.93$ $87.65\pm 2.24$
10 D3RE VI $100\%$ $0.49\pm 0.39$ $3.27\pm 2.00$ $20.30\pm 0.55$ $42.65\pm 26.87$ $102.01\pm 2.43$
10 ISA-DRE (ours) VI $0\%$ $0.94\pm 1.26$ $7.59\pm 0.54$ $\mathbf{17.92}\pm\mathbf{1.15}$ $18.45\pm 1.92$ $139.62\pm 7.72$
10 ISA-DRE (ours) VI $50\%$ $0.70\pm 1.24$ $7.59\pm 0.83$ $17.93\pm 0.84$ $19.11\pm 1.39$ $128.78\pm 2.91$
10 ISA-DRE (ours) VI $100\%$ $\mathbf{-1.97}\pm\mathbf{0.40}$ $3.98\pm 8.42$ $18.89\pm 2.61$ $29.40\pm 16.32$ $55.54\pm 8.25$
10 ISA-DRE (ours) VI CIA $1.17\pm 0.08$ $\mathbf{-9.66}\pm\mathbf{0.01}$ $18.23\pm 0.36$ $\mathbf{17.29}\pm\mathbf{0.48}$ $\mathbf{-162.14}\pm\mathbf{3.98}$
10 ISA-DRE (ours) Uni. $0\%$ $0.69\pm 0.30$ $8.53\pm 1.25$ $17.66\pm 0.60$ $13.19\pm 0.47$ $\mathbf{-160.03}\pm\mathbf{10.48}$
10 ISA-DRE (ours) Uni. $50\%$ $\mathbf{-0.92}\pm\mathbf{0.59}$ $8.15\pm 1.11$ $\mathbf{17.62}\pm\mathbf{0.70}$ $13.73\pm 0.13$ $155.48\pm 16.12$
10 ISA-DRE (ours) Uni. $100\%$ $0.52\pm 0.39$ $\mathbf{-8.82}\pm\mathbf{1.32}$ $18.41\pm 0.01$ $51.34\pm 51.17$ $110.97\pm 33.15$
10 ISA-DRE (ours) Uni. CIA $0.80\pm 0.15$ $7.34\pm 0.14$ $17.70\pm 0.77$ $\mathbf{13.04}\pm\mathbf{0.96}$ $154.75\pm 2.27$
10 ISA-DRE (ours) LN $0\%$ $\mathbf{-0.92}\pm\mathbf{0.19}$ $6.28\pm 0.91$ $\mathbf{18.46}\pm\mathbf{0.00}$ $13.47\pm 1.47$ $131.71\pm 2.71$
10 ISA-DRE (ours) LN $50\%$ $0.58\pm 0.47$ $6.51\pm 0.19$ $18.86\pm 0.63$ $13.32\pm 0.84$ $129.67\pm 1.89$
10 ISA-DRE (ours) LN $100\%$ $0.38\pm 0.21$ $57.55\pm 25.63$ $38.45\pm 24.31$ $22.08\pm 4.13$ $29.70\pm 14.58$
10 ISA-DRE (ours) LN CIA $0.65\pm 0.45$ $\mathbf{-8.53}\pm\mathbf{0.21}$ $18.48\pm 0.98$ $\mathbf{12.98}\pm\mathbf{0.45}$ $\mathbf{-148.53}\pm\mathbf{3.16}$
Table 1: Density estimation results on five real-world tabular datasets with complex, non-Gaussian structures. Values indicate negative log-likelihood (NLL; lower is better), reported as $\text{mean}\pm\text{std}$ over 3 runs. Results are shown across varying function evaluations ($\text{NFE}\in\{2,5,10,50\}$), time samplers (TS; see Sec. 4.3), and secant-tangent supervisions (STS; see Sec. 4.3). Bold entries mark the best mean NLL for each NFE-TS-dataset setting. Additional results (NFE $\in\{2,5,50\}$) are provided in Tab. 4.
Figure 4: Density estimation performance with a fixed number of function evaluations (NFE =2=2) on structured and multi-modal datasets. The ISA-DRE (ours) markedly outperforms DRE-$\infty$ and D3RE in low-NFE regimes
Real-world Tabular Datasets.
We evaluate ISA-DRE on five challenging tabular datasets with complex, non-Gaussian structures (Grathwohl et al. 2018). As shown in Tabs. 1 and 4, ISA-DRE achieves state-of-the-art density estimation, especially under low-NFE settings. At $\text{NFE}=2$, it outperforms prior methods, e.g., on BSDS300, 234.21-234.21 vs. D3RE’s 149.53-149.53; on MINIBOONE, $20.05$ vs. DRE-$\infty$’s $41.55$. This advantage persists with more compute: at $\text{NFE}=50$, VI+CIA achieves best-in-class results on GAS (8.19-8.19) and BSDS300 (150.54-150.54), outperforming baselines by over $25\%$. These results highlight ISA-DRE’s robustness and generalization, with the VI+CIA setup particularly effective at modeling complex dependencies.
5.3 Mutual information Estimation
Mutual information (MI) quantifies the statistical dependence between two random variables $\mathbf{X}\sim p(\mathbf{x})$ and $\mathbf{Y}\sim q(\mathbf{y})$, and is defined as $\text{MI}(\mathbf{X},\mathbf{Y})=\mathbb{E}_{p(\mathbf{x},\mathbf{y})}\left[\log\frac{p(\mathbf{x},\mathbf{y})}{p(\mathbf{x})q(\mathbf{y})}\right]$. Since this formulation involves the expectation of a log density ratio, MI estimation naturally reduces to a DRE problem. To evaluate the robustness of our method, we introduce two challenging MI estimation benchmarks: (1) geometrically pathological distributions, where $(\mathbf{X},\mathbf{Y})$ lie on complex non-linear manifolds; and (2) high‑discrepancy Gaussian pairs, with strong correlations and anisotropic covariance structures.
Geometrically Pathological Distributions.
We evaluate ISA-DRE on four geometrically pathological distributions from Czyż et al. (2023): Asinh Mapping, Additive Noise, Half-Cube Map, and Edge-singular Gauss (Fig. 5). On the first two (Asinh Mapping and Additive Noise), ISA-DRE achieves the lowest MSE ($0.0010$ and $0.0031$), significantly outperforming D3RE and DRE-$\infty$, the latter of which fails entirely (MSEs of $3.2582$ and $39.8936$). On Edge-singular Gauss, DRE-$\infty$ performs best ($0.0005$), slightly ahead of D3RE and ISA-DRE. On Half-Cube Map, D3RE attains the lowest error ($0.0018$), while ISA-DRE remains competitive ($0.0062$), both far exceeding DRE-$\infty$ ($6.9134$). Overall, DRE-$\infty$ struggles on these complex benchmarks, while ISA-DRE and D3RE exhibit strong robustness. ISA-DRE achieves the best or comparable results in most cases.
(a) Asinh Mapping
(b) Additive Noise
(c) Half-Cube Map
(d) Edge-singular Gauss
Figure 5: Comparison of ISA-DRE (ours) with DRE-$\infty$ and D3RE on four geometrically pathological distributions. Mean Squared Error (MSE) values for each method are reported. ISA-DRE achieves the best or comparable results.
$d=40$ (MI =10=10) $d=80$ (MI =20=20) $d=120$ (MI =30=30) $d=160$ (MI =40=40)
NFE Method TS STS Est. MI MSE Est. MI MSE Est. MI MSE Est. MI MSE
$10$ DRE-$\infty$ VI $100\%$ $9.48\pm 0.06$ $0.27$ $19.27\pm 0.04$ $0.54$ $28.37\pm 0.05$ $2.66$ $37.34\pm 0.06$ $7.08$
$10$ D3RE VI $100\%$ $10.13\pm 0.04$ $\mathbf{0.02}$ $20.45\pm 0.03$ $\mathbf{0.21}$ $27.22\pm 0.03$ $7.72$ $32.27\pm 0.04$ $59.70$
$10$ ISA-DRE (ours) VI $0\%$ $10.45\pm 0.04$ $0.20$ $21.14\pm 0.04$ $1.29$ $30.86\pm 0.03$ $0.75$ $41.64\pm 0.03$ $2.68$
$10$ ISA-DRE (ours) VI $50\%$ $10.70\pm 0.04$ $0.49$ $21.39\pm 0.03$ $1.94$ $32.00\pm 0.04$ $3.99$ $42.19\pm 0.03$ $4.79$
$10$ ISA-DRE (ours) VI $100\%$ $5.90\pm 0.03$ $16.82$ $15.26\pm 0.03$ $22.45$ $73.87\pm 0.03$ $1924.57$ $37.11\pm 0.04$ $8.33$
$10$ ISA-DRE (ours) VI CIA $1.09\pm 0.05 1.1$91.19 $20.92\pm 0.03$ $0.86$ $30.22\pm 0.04$ $\mathbf{0.05}$ $40.74\pm 0.04$ $\mathbf{0.54}$
$10$ ISA-DRE (ours) Uni. $0\%$ $10.68\pm 0.06$ $0.47$ $21.76\pm 0.05$ $3.09$ $30.49\pm 0.04$ $0.25$ $41.00\pm 0.02$ $0.99$
$10$ ISA-DRE (ours) Uni. $50\%$ $10.61\pm 0.05$ $\mathbf{0.37}$ $22.26\pm 0.05$ $5.12$ $30.28\pm 0.04$ $0.08$ $40.33\pm 0.02$ $\mathbf{0.11}$
$10$ ISA-DRE (ours) Uni. $100\%$ $12.30\pm 0.06$ $5.31$ $31.83\pm 0.05$ $139.89$ $102.81\pm 0.04$ $5300.98$ $46.62\pm 0.10$ $43.86$
$10$ ISA-DRE (ours) Uni. CIA $11.40\pm 0.06$ $1.96$ $21.46\pm 0.04$ $\mathbf{2.12}$ $30.24\pm 0.05$ $\mathbf{0.06}$ $39.48\pm 0.03$ $0.27$
$10$ ISA-DRE (ours) LN $0\%$ $11.49\pm 0.05$ $2.23$ $22.09\pm 0.03$ $4.38$ $32.56\pm 0.03$ $6.57$ $39.84\pm 0.03$ $0.03$
$10$ ISA-DRE (ours) LN $50\%$ $11.62\pm 0.05$ $2.61$ $21.97\pm 0.03$ $3.88$ $31.33\pm 0.02$ $1.77$ $38.68\pm 0.03$ $1.76$
$10$ ISA-DRE (ours) LN $100\%$ $26.91\pm 0.04$ $286.02$ $55.68\pm 0.04$ $1273.42$ $39.60\pm 0.02$ $92.13$ $32.71\pm 0.03$ $53.08$
$10$ ISA-DRE (ours) LN CIA $10.84\pm 0.05$ $\mathbf{0.71}$ $20.33\pm 0.04$ $\mathbf{0.11}$ $30.71\pm 0.05$ $\mathbf{0.51}$ $39.89\pm 0.07$ $\mathbf{0.02}$
Table 2: Mutual information estimation under high-discrepancy settings (MI $\in\{10,20,30,40\}$ nats). We report the estimated mutual information (mean $\pm$ std over 5 seeds) and MSE across different numbers of function evaluations (NFE $\in\{2,5,10,50\}$), time samplers (TS; see Sec. 4.3), and secant-tangent supervisions (STS; see Sec. 4.3). Bolded MSE values indicate the best performance for each NFE-TS-$d$ combination. Additional results for NFE $\in\{2,5,50\}$ are provided in Tab. 5.
High-dimensional & High-discrepancy Distributions.
We evaluate ISA-DRE on blockwise correlated Gaussians designed to induce extreme high-discrepancy (MI $\geq 20$ nats), where traditional estimators suffer from the density-chasm problem (Rhodes, Xu, and Gutmann 2020). Following Choi et al. (2022), we implement ISA-DRE and summarize the results in Tab. 5. ISA-DRE consistently succeeds where others collapse. At MI =40=40 ($d=160$) with only 2 function evaluations (NFE =2=2), ISA-DRE achieves near-perfect estimation (MSE =0.72=0.72), while DRE-$\infty$ and D3RE fail catastrophically (MSE =1215.69=1215.69 and $500.04$, respectively). With more computation (NFE =50=50), ISA-DRE (VI+CIA) reduces error to near-zero (MSE =0.01=0.01), outperforming all baselines by 3+ orders of magnitude. Notably, ISA-DRE maintains sub-1.0 MSE across all tested dimensions and NFEs at MI =40=40, confirming its robustness in extreme regimes. These results highlight the strength of ISA-DRE in mitigating density-chasm problem that fundamentally break existing estimators.
5.4 Ablation Study
Our ablation studies isolate two key components driving ISA-DRE’s stability in high-discrepancy settings: VI sampling and CIA supervision. Together, they enable stable and accurate estimation where prior methods fail.
Ablation: Time Sampler.
VI sampling consistently outperforms LN and Uniform strategies, especially at low NFEs. (1) In MI estimation (see Tabs. 2 and 5), VI sampling shows superior accuracy: at MI =40=40 and $\text{NFE}=2$, VI achieves a MSE of $0.72$, substantially lower than LN ($22.16$) and Uniform ($181.41$). This advantage persists with increased computation: at $\text{NFE}=5$, VI still outperforms LN ($0.30$ vs. $6.78$). By contrast, Uniform sampling remains unstable even at $\text{NFE}=50$, with MSE exceeding $600$. (2) In density estimation (see Tabs. 1 and 4), VI sampling consistently yields the best performance, especially on high-dimensional datasets. At $\text{NFE}=2$, VI improves the negative log-likelihood (NLL) by over $50\%$ on GAS (14.95-14.95 vs. 10.5-10.5 for Uniform) and by over $35\%$ on POWER (5.07-5.07 vs. 1.82-1.82). While LN sampling performs competitively on low-dimensional data (e.g., BSDS300 with 211.86-211.86 at $\text{NFE}=2$), it suffers from instability in higher dimensions (e.g., GAS with $29.41$ at $\text{NFE}=5$). Uniform sampling shows strong dataset sensitivity, performing reasonably on MINIBOONE ($18.73$ at $\text{NFE}=2$), but failing catastrophically on BSDS300 (27.3-27.3 at $\text{NFE}=2$). More results are reported in Fig. 6.
Ablation: Secant-Tangent Supervision (STS).
Among the tested configurations, CIA consistently offers the best stability and performance. (1) For MI estimation (see Tabs. 2 and 5), secant-only supervision ($0\%$ STS) is remarkably robust: at MI =40=40 with $\text{NFE}=2$, it achieves MSE =0.72=0.72, a $77\times$ improvement over $100\%$ STS ($55.97$). CIA further improves accuracy at higher NFEs: at $d=160$, $\text{NFE}=50$, CIA with VI sampling achieves near-zero error (MSE =0.01=0.01). In contrast, tangent-only supervision ($100\%$ STS) is often unstable and can catastrophically fail, as seen at $d=120$, $\text{NFE}=2$ (MSE =2804.96=2804.96). CIA also scales better with dimensionality: at $d=120$, MI =30=30, it reduces MSE from $2804.96$ to $3.21$ under VI sampling. (2) In density estimation (see Tabs. 1 and 4), CIA proves especially effective at higher NFEs, achieving the best performance on $80\%$ of datasets when $\text{NFE}\geq 10$ (e.g., MINIBOONE NLL =12.58=12.58 at $\text{NFE}=50$). Full supervision remains unstable, particularly for high-dimensional data, for instance, on HEPMASS at $\text{NFE}=5$, $100\%$ STS degrades performance by over 100% compared to $0\%$ STS (NLL =41.06=41.06 vs. $15.87$). CIA mitigates this instability while retaining flexibility: on POWER at $\text{NFE}=2$, it achieves the best NLL (5.07-5.07), outperforming $100\%$ STS by over $40\%$. Tangent-only ($0\%$ STS) performs best in low-NFE settings, leading on 60% of datasets at $\text{NFE}\leq 5$ (e.g., BSDS300 234.21-234.21 at $\text{NFE}=2$). These results confirm that CIA provides the best balance between stability and expressiveness, making it well-suited for real-world high-discrepancy tasks.
6 Conclusion and Future Work
We proposed ISA-DRE (Interval-annealed Secant Alignment Density-Ratio Estimation), a novel framework that addresses the key trade-off in modern density ratio estimation (DRE) between accuracy and efficiency. Unlike existing tangent-based methods that rely on expensive numerical integration during inference, ISA-DRE directly learns the secant: the expectation of tangents over an interval. We further provide a theoretical guarantee that secants have lower variance than tangents, making them more suitable for neural approximation (Proposition 4.1). At its core is the Secant Alignment Identity, a consistency condition enabling estimating the desired density ratio in any step without sacrificing accuracy. To ensure stable and consistent training, we propose Contraction Interval Annealing, a curriculum strategy that progressively refines the learning interval to enhance convergence. ISA-DRE achieves competitive performance in low-NFE regimes, making it particularly well-suited for real-time or interactive applications. Looking ahead, ISA-DRE presents challenges such as the need for higher-capacity models due to its richer input $(\mathbf{x},l,t)$, and sensitivity to the choice of interpolant—an open problem shared by prior tangent-based DRE methods.
Appendix A Proofs
A.1 Preliminary Lemmas and Its Proof
Lemma A.1.
Consider the auxiliary random variable $\eta$ generated through the following two-step sampling procedure: (1): sample $l$ and $t$ independently from the distribution with density $p(\cdot)$ on $[0,1]$, conditioned on $l\leq t$; (2): given $l$ and $t$, sample $\eta$ uniformly from the interval $[l,t]$, denoted as $\eta\mid(l,t)\sim\mathcal{U}[l,t]$. Then $\eta$ possesses the marginal probability density function $p(\cdot)$ on $[0,1]$.
Proof.
The marginal probability $P(\eta\in A)$ for any Borel set $A\subseteq[0,1]$ is given by:
$P(\eta\in A)$ (22)
$=$ $\frac{1}{P(l\leq t)}\iint\limits_{\begin{subarray}{c}0\leq l\leq t\leq 1\end{subarray}}P(\eta\in A\mid l,t)p(l)p(t)\mathrm{d}l\mathrm{d}t$
$=$ $\frac{1}{P(l\leq t)}\iint\limits_{l\leq t}\left(\frac{1}{t-l}\int_{A}\mathbf{1}_{[l,t]}(\tau)d\tau\right)p(l)p(t)\mathrm{d}l\mathrm{d}t$
$=$ $\frac{1}{P(l\leq t)}\int_{A}\left(\iint\limits_{\begin{subarray}{c}l\leq\tau\leq t\end{subarray}}\frac{p(l)p(t)}{t-l}\mathrm{d}l\mathrm{d}t\right)\mathrm{d}\tau.$
The inner double integral simplifies via Fubini’s theorem and the definition of the joint density:
$\iint\limits_{\begin{subarray}{c}l\leq\tau\leq t\end{subarray}}\frac{p(l)p(t)}{t-l}\mathrm{d}l\mathrm{d}t$ $=\int_{0}^{\tau}\int_{\tau}^{1}\frac{p(l)p(t)}{t-l}\mathrm{d}t\mathrm{d}l$ (23)
$=p(\tau)\cdot P(l\leq t),$
where the last equality follows from the observation that the integral represents the joint density normalization.
Substituting this result yields:
$P(\eta\in A)$ $=\frac{1}{P(l\leq t)}\int_{A}p(\tau)\cdot P(l\leq t)d\tau$ (24)
$=\int_{A}p(\tau)d\tau,$
confirming that $\eta$ indeed has marginal density $p(\cdot)$. ∎
Proposition A.2.
Assume the score function $s_{t}(\mathbf{x},\tau)$ is $\lambda$-Lipschitz continuous in $\tau$. Then for fixed $\mathbf{x}$ and $l\in[0,1)$, the secant function $u(\mathbf{x},l,t)$ is $\frac{\lambda}{2}$-Lipschitz continuous in $t$:
$|u(\mathbf{x},l,t_{1})-u(\mathbf{x},l,t_{2})|\leq\frac{\lambda}{2}|t_{1}-t_{2}|,\forall t_{1},t_{2}\in(l,1]$ (25)
Proof.
Fix $\mathbf{x}$ and $l$, and consider $t_{1},t_{2}>l$. Define the auxiliary function:
$g(t)=\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau,$ (26)
so that the secant function can be expressed as:
$u(\mathbf{x},l,t)=\frac{g(t)}{t-l}.$ (27)
The derivative of $u$ with respect to $t$ is:
$\frac{\mathrm{d}u}{\mathrm{d}t}=\frac{g^{\prime}(t)(t-l)-g(t)}{(t-l)^{2}}.$ (28)
By the Fundamental Theorem of Calculus, $g^{\prime}(t)=s_{t}(\mathbf{x},t)$, thus the absolute value of numerator, i.e., $g^{\prime}(t)(t-l)-g(t)$, satisfies:
$\quad\left|g^{\prime}(t)(t-l)-g(t)\right|$ (29)
$=\left|s_{t}(\mathbf{x},t)(t-l)-\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau\right|$
$=\left|\int_{l}^{t}\left[s_{t}(\mathbf{x},t)-s_{t}(\mathbf{x},\tau)\right]\mathrm{d}\tau\right|$
$\leq\int_{l}^{t}\left|s_{t}(\mathbf{x},t)-s_{t}(\mathbf{x},\tau)\right|\mathrm{d}\tau$
$\leq\int_{l}^{t}\lambda|t-\tau|\mathrm{d}\tau\quad(\star)$
$=\lambda\left(\frac{1}{2}t^{2}-tl+\frac{1}{2}l^{2}\right)$
$=\frac{\lambda}{2}(t-l)^{2},$
where $(\star)$ holds because the score function $s_{t}(\mathbf{x},\tau)$ is $\lambda$-Lipschitz continuous in $\tau$, which leads to $|s_{t}(\mathbf{x},t)-s_{t}(\mathbf{x},\tau)|\leq\lambda|t-\tau|,\forall\tau\in[l,t]$.
Substituting this bound into Eq. 28:
$\left|\frac{\mathrm{d}u}{\mathrm{d}t}\right|\leq\frac{\frac{\lambda}{2}(t-l)^{2}}{(t-l)^{2}}=\frac{\lambda}{2}.$ (30)
By the Mean Value Theorem, for any $t_{1},t_{2}>l$:
$|u(\mathbf{x},l,t_{2})-u(\mathbf{x},l,t_{1})|$ $\leq\sup_{\xi\in[t_{1},t_{2}]}\left|\frac{\partial u}{\partial t}(\xi)\right|\cdot|t_{2}-t_{1}|$ (31)
$\leq\frac{\lambda}{2}|t_{2}-t_{1}|,$
which completes the proof. ∎
A.2 Proof of Proposition 4.1
Proof.
To establish the variance inequality, we introduce an auxiliary random variable $\eta$ whose conditional distribution given $l$ and $t$ is uniform on $[l,t]$, denoted as $\eta\mid(l,t)\sim\mathcal{U}[l,t]$. Define the composite random variable $Z=s_{t}(\mathbf{x},\eta)$.
We now apply the law of total variance to $S$, conditioning on the $\sigma$-algebra generated by $(l,t)$. The decomposition yields:
$\text{Var}_{p(\tau)}(S)=$ $\mathbb{E}_{p(l,t)}\left[\text{Var}_{\mathcal{U}[l,t]}\left(Z\mid l,t\right)\right]$ (32)
$+\text{Var}_{p(l,t)}\left(\mathbb{E}_{\mathcal{U}[l,t]}\left[Z\mid l,t\right]\right),$
where all conditional expectations and variances are computed with respect to the uniform measure $\mathcal{U}[l,t]$ on $\eta$.
The second term on the right hand side of Eq. 32 is equal to the variance of the secant function:
$\mathbb{E}_{\mathcal{U}[l,t]}\left[Z\mid l,t\right]$ $=\mathbb{E}_{\eta\sim\mathcal{U}[l,t]}\left[s_{t}(\mathbf{x},\eta)\mid l,t\right]$ (33)
$=\frac{1}{t-l}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau$
$=u(\mathbf{x},l,t)=U.$
Substituting this identity into Eq. 32:
$\text{Var}_{p(\tau)}(S)=$ $\mathbb{E}_{p(l,t)}\left[\text{Var}_{\mathcal{U}[l,t]}\left(Z\mid l,t\right)\right]$ (34)
$+\text{Var}_{p(l,t)}(U).$
For the first term on the right hand side of Eq. 32, the conditional variance term is non-negative:
$\text{Var}_{\mathcal{U}[l,t]}\left(Z\mid l,t\right)$ $=\text{Var}_{\eta\sim\mathcal{U}[l,t]}\left(s_{t}(\mathbf{x},\eta)\mid l,t\right)$ (35)
$=\mathbb{E}_{\mathcal{U}[l,t]}\left[\left|s_{t}(\mathbf{x},\eta)-U\right|^{2}\mid l,t\right]$
$\geq 0.$
which implies:
$\text{Var}_{p(\tau)}(S)\geq\text{Var}_{p(l,t)}(U).$ (36)
Equality holds if and only if for $p$-almost every pair $(l,t)$:
$\text{Var}_{\eta\sim\mathcal{U}[l,t]}\left(s_{t}(\mathbf{x},\eta)\mid l,t\right)=0,$ (37)
which necessitates $s_{t}(\mathbf{x},\cdot)$ being constant Lebesgue-almost everywhere on $[l,t]$. As the intervals $[l,t]$ densely cover $[0,1]$ when $(l,t)$ varies within the support of $p(\cdot)$, and since Lemma A.1 establishes that $\eta$ has marginal density $p(\tau)$, this implies $s_{t}(\mathbf{x},\tau)$ must be constant for $p$-almost every $\tau\in[0,1]$.
A.3 Proof of Proposition 4.2
Proof.
The proposition is a biconditional statement, which requires us to prove two directions.
First, we prove the forward direction ($\Rightarrow$).
We assume that $u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)$ is identical to the true secant function and demonstrate that it must satisfy both the boundary and consistency conditions. Assume
$u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)=\frac{1}{t-l}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau.$ (38)
We begin by verifying the boundary condition. The limit of $u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},t_{0},t)$ as $t\to t_{0}$ takes the indeterminate form $\frac{0}{0}$. We can therefore apply L’Hôpital’s Rule:
$\lim_{t\to t_{0}}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},t_{0},t)$ $=\lim_{t\to t_{0}}\frac{\int_{t_{0}}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau}{t-t_{0}}$ (39)
$=\lim_{t\to t_{0}}\frac{\frac{\partial}{\partial t}\left(\int_{t_{0}}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau\right)}{\frac{\partial}{\partial t}(t-t_{0})}$
$=\lim_{t\to t_{0}}\frac{s_{t}(\mathbf{x},t)}{1}=s_{t}(\mathbf{x},t_{0}).$
This confirms that the boundary condition is satisfied.
Next, we verify the consistency condition. We must show that the expression $u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)+(t-l)\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)$ simplifies to $s_{t}(\mathbf{x},t)$. To do this, we first compute the partial derivative $\frac{\mathrm{d}u^{\boldsymbol{\theta}^{\star}}}{\mathrm{d}t}$ using the product rule for differentiation on the expression $u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)=\frac{1}{t-l}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau$:
$\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)$ (40)
$=$ $\frac{\mathrm{d}(t-l)^{-1}}{\mathrm{d}t}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau+\frac{1}{t-l}\frac{\mathrm{d}}{\mathrm{d}t}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau$
$=$ $\frac{-1}{(t-l)^{2}}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau+\frac{1}{t-l}s_{t}(\mathbf{x},t)$
$=$ $\frac{-1}{(t-l)^{2}}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau+\frac{s_{t}(\mathbf{x},t)}{t-l}.$
Now, substituting this derivative back into the consistency condition expression yields:
$u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)+(t-l)\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)$ (41)
$=$ $\frac{1}{t-l}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau-\frac{1}{t-l}\int_{l}^{t}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau+s_{t}(\mathbf{x},t)$
$=$ $s_{t}(\mathbf{x},t).$
The consistency condition is therefore also satisfied. This completes the proof of the forward direction.
Finnaly, we now prove the reverse direction ($\Leftarrow$).
We assume that a function $u^{\boldsymbol{\theta}^{\star}}$ satisfies both the boundary and consistency conditions. Our objective is to prove that $u^{\boldsymbol{\theta}^{\star}}$ must be uniquely determined and equal to the true secant function. The consistency condition itself is a differential equation that governs the evolution of $u^{\boldsymbol{\theta}^{\star}}$ with respect to time $t$. For $t\neq l$, we can rearrange the equation into the standard form of a first-order linear ordinary differential equation:
$\frac{\mathrm{d}}{\mathrm{d}t}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)+\frac{1}{t-l}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)=\frac{1}{t-l}s_{t}(\mathbf{x},t).$ (42)
This type of differential equation can be solved using the method of integrating factors. The integrating factor, denoted $I(t)$, is given by:
$I(t)=\exp\left\{\int\frac{1}{t-l}\mathrm{d}t\right\}=\exp(\ln|t-l|)=|t-l|.$ (43)
Without loss of generality, let us consider the case where $t>l$, so the integrating factor is (tl)(t-l). Multiplying the standard-form ODE by this factor yields:
$(t-l)\frac{\mathrm{d}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)}{\mathrm{d}t}+u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)$ $=s_{t}(\mathbf{x},t)$ (44)
$\Rightarrow\frac{\mathrm{d}}{\mathrm{d}t}\left[(t-l)u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t)\right]$ $=s_{t}(\mathbf{x},t).$
We can now integrate both sides with respect to $\tau$ from the initial point $l$ to a generic endpoint $t^{\prime}$:
$\int_{l}^{t^{\prime}}\frac{\mathrm{d}}{\mathrm{d}\tau}\left[(\tau-l)u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,\tau)\right]\mathrm{d}\tau=\int_{l}^{t^{\prime}}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau.$ (45)
Applying the Fundamental Theorem of Calculus to the left-hand side gives:
$\left[(\tau-l)u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,\tau)\right]_{\tau=l}^{\tau=t^{\prime}}=\int_{l}^{t^{\prime}}s_{t}(\mathbf{x},\tau)\mathrm{d}\tau.$ (46)
Evaluating the expression at the bounds, we have:
$\int_{l}^{t^{\prime}}s_{\tau}(\mathbf{x},\tau)\mathrm{d}\tau$ (47)
$=$ $(t^{\prime}-l)u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t^{\prime})-\lim_{\tau\to l^{+}}(\tau-l)u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,\tau)$
$=$ $(t^{\prime}-l)u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t^{\prime})-\lim_{\tau\to l^{+}}(\tau-l)\cdot\lim_{\tau\to l^{+}}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,\tau)$
$=$ $(t^{\prime}-l)u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t^{\prime})-0\cdot s_{t}(\mathbf{x},l)\quad(\star)$
$=$ $(t^{\prime}-l)u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t^{\prime}).$
Here, the equality $(\star)$ holds according to the boundary condition. We know that $\lim_{\tau\to l^{+}}u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,\tau)$ converges to the finite value $s_{l}(\mathbf{x},l)$.
The limit term vanishes, leaving us with a direct algebraic relationship:
$(t^{\prime}-l)u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t^{\prime})=\int_{l}^{t^{\prime}}s_{\tau}(\mathbf{x},\tau)\mathrm{d}\tau.$ (48)
For any $t^{\prime}\neq l$, we can divide by $(t^{\prime}-l)$ to solve for $u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t^{\prime})$:
$u^{\boldsymbol{\theta}^{\star}}(\mathbf{x},l,t^{\prime})=\frac{1}{t^{\prime}-l}\int_{l}^{t^{\prime}}s_{\tau}(\mathbf{x},\tau)\mathrm{d}\tau.$ (49)
This is precisely the definition of the true secant function. Furthermore, the Picard-Lindelöf theorem guarantees that the solution to this linear initial value problem is unique. Thus, any function $u^{\boldsymbol{\theta}^{\star}}$ satisfying the consistency and boundary conditions must be the true secant function. This completes the proof of the reverse direction. ∎
Appendix B Experimental Settings and Results for Density Estimation
B.1 Training Procedure
In each training step, we sample a batch of pairs $(\mathbf{x}_{0},\mathbf{x}_{1})$ from the source and target distributions $p_{0}\times p_{1}$, respectively. We also sample a time $t$ using a time sampler $p(\cdot)$ defined over $[0,1]$. The interpolated sample $\mathbf{x}_{t}$ is then constructed via $\mathbf{x}_{t}=a_{t}\mathbf{x}_{0}+b_{t}\mathbf{x}_{1}$. We use the coefficients $(a_{t},b_{t})=(1-t,t)$ for tabular datasets and $(a_{t},b_{t})$ corresponding to the VPSDE (Song et al. 2021) schedule for other datasets, following the setup in Choi et al. (2022); Chen et al. (2025).
B.2 Structured and Multi-modal Distributions
Datasets.
Our model’s performance is evaluated on a comprehensive suite of nine 2D synthetic datasets, including swissroll, circles, rings, moons, 8gaussians, pinwheel, 2spirals, checkerboard and tree. These benchmarks are specifically chosen to probe the model’s capacity to learn distributions with diverse and challenging characteristics, ranging from complex topologies to discontinuous densities. Eight of these are standard benchmarks (Chen et al. 2025) selected to test the model on distributions with multi-modal (8gaussians, moons), disconnected (circles, rings), intricately structured (swissroll, pinwheel, 2spirals), and discontinuous (checkerboard) properties. These datasets are known to be challenging for generative models due to their complex geometric and topological features. Additionally, we use the tree dataset from Bansal, Gee, and Fletcher (2023), which is specifically designed to assess a model’s ability to generate sharp, branching topological structures. Together, these datasets form a comprehensive testbed for evaluating the performance of our model.
B.3 Real-world Tabular Distributions
Datasets.
We also assess the scalability and generalizability of our model on five high-dimensional, real-world tabular datasets. These datasets, originating from diverse scientific and industrial domains such as high-energy physics, and power grid monitoring, are standard benchmarks for tabular data modeling (Grathwohl et al. 2018). Unlike the synthetic examples, the underlying generative processes of these datasets are unknown and they exhibit complex, non-Gaussian correlations and potentially noisy features. Therefore, they provide a crucial testbed for evaluating a model’s ability to capture the intricate data distributions encountered in practical applications. We use the same pre-processing and data splits as established in Grathwohl et al. (2018) to ensure a fair comparison. The key characteristics of these datasets are summarized in Tab. 3.
Table 3: Summary of real-world tabular datasets used in our experiments. “Dimension” refers to the feature dimensionality and “Samples” denotes the total number of instances in each dataset.
Dataset Dimension Samples Batch Size
POWER 6 1,659,917 50,000
GAS 8 852,174 40,000
HEPMASS 21 315,123 20,000
MINIBOONE 43 29,556 1,000
BSDS300 63 1,000,000 50,000
B.4 Ablation Studies
Ablation Studies on Time Samplers.
Ablations in Fig. 6 compare Uniform, LN, and VI samplers at $\text{NFE}=2$ and $\text{NFE}=10$. At $\text{NFE}=2$, all samplers learn the overall support but with minor differences in mode sharpness. Increasing to $\text{NFE}=10$, the VI sampler markedly sharpens density estimates and reduces bias across datasets, whereas Uniform and LN show slower gains, highlighting the importance of variance‑based time sampling for high‑fidelity density estimation.
(a) $\text{NFE}=2$
(b) $\text{NFE}=10$
Figure 6: Ablation studies on time samplers for density estimation with different NFE settings. Experiments are done on various structured and multi-modal datasets.
Appendix C Experimental Settings and Results for Mutual Information Estimation
C.1 Geometrically Pathological Distributions
C.2 Training Procedure
In each training step, we sample a batch of pairs $(\mathbf{x}_{0},\mathbf{x}_{1})$ from the source and target distributions, $p_{0}$ and $p_{1}$, respectively. We also sample a time $t$ using a time sampler $p(\cdot)$ defined over $[0,1]$. The interpolated sample $\mathbf{x}_{t}$ is then constructed via $\mathbf{x}_{t}=a_{t}\mathbf{x}_{0}+b_{t}\mathbf{x}_{1}$. Following standard practice in score-based modeling, we use the coefficients $(a_{t},b_{t})$ corresponding to the VPSDE schedule following Song et al. (2021); Choi et al. (2022); Chen et al. (2025).
Dataset.
We validate our method on four geometrically pathological distributions proposed in Czyż et al. (2023), including Asinh Mapping, Additive Noise, Half-Cube Map and Edge-singular Gauss.
  • Asinh Mapping: $\left(\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\right)=\left(\text{asinh}(\mathbf{X}),\text{asinh}(\mathbf{Y})\right)$, where $\text{asinh}(\mathbf{x})=\log\left(\mathbf{x}+\sqrt{\mathbf{x}^{2}+1}\right)$ denotes the inverse hyperbolic sine. This transformation generates peak-concentrated densities, compressing distribution tails into high-curvature central regions that challenge kernel-based estimators and induce gradient instability.
  • Additive Noise: Let $\mathbf{X}\sim\mathcal{U}(0,1)$, $\mathbf{N}\sim\mathcal{U}(-\epsilon,\epsilon)$ (independent), $\mathbf{Y}=\mathbf{X}+\mathbf{N}$. The MI between $\mathbf{X}$ and $\mathbf{Y}$ is defined as $I(\mathbf{X};\mathbf{Y})=\log(2\epsilon)+0.5\ (\epsilon\leq 0.5)$. The boundary-fragmented discontinuities and piecewise-constant densities sabotage differentiable estimators through support-set irregularities.
  • Half-Cube Map: Let $\text{half-cube}(\mathbf{x})=|\mathbf{x}|^{3/2}\text{sign}(\mathbf{x})$ be a homeomorphism and apply it to Gaussian variables $\mathbf{X}$ and $\mathbf{Y}$: $\left(\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\right)=\left(\text{half-cube}(\mathbf{X}),\text{half-cube}(\mathbf{Y})\right)$. While $I(\mathbf{X}^{\prime};\mathbf{Y}^{\prime})=I(\mathbf{X};\mathbf{Y})$, the heavy-tail distortion from cubic mapping disrupts local density estimation in non-parametric methods.
  • Edge-singular Gauss: Consider Gaussian variables $(\mathbf{X},\mathbf{Y})\sim\mathcal{N}\left(\mathbf{0},[[1,\rho],[\rho,1]]\right)$ with $\rho$ being the correlation coefficient between $\mathbf{X}$ and $\mathbf{Y}$. Their MI can be evaluated by $I(\mathbf{X};\mathbf{Y})=-0.5\log(1-\rho^{2})$. Exhibits edge-singular characteristics where MI estimation becomes numerically unstable as $\rho\to 0^{+}$ (vanishing gradients) or $\rho\to 1^{-}$ (divergence).
C.3 High-dimensional & High-discrepancy Distributions
Datasets.
To evaluate model performance under challenging conditions, we construct a scenario with high discrepancy between the source and target distributions, a known trigger for the density-chasm problem (Rhodes, Xu, and Gutmann 2020). This problem arises when the probability path between two distributions passes through a region of near-zero density, causing DRE-based models to fail. Our experiment is designed to test whether ISA-DRE can effectively bridge this chasm. Specifically, the covariance matrix is block diagonal, with each $2\times 2$ block defined as $\Lambda=\left[\left[1,\rho\right],\left[\rho,1\right]\right]$, inducing strong pairwise correlations within each block and maintaining independence across blocks. This structure introduces localized dependencies and global sparsity, resulting in a highly ill-conditioned and low-rank covariance matrix. Such statistical characteristics pose significant challenges for density ratio estimation (Choi et al. 2022). We define the source and target distributions as $q_{0}(\mathbf{x})=\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ and $q_{1}(\mathbf{x})=\mathcal{N}(\mathbf{0},\Sigma)$, where $\Sigma$ is composed of repeated $\Lambda$ blocks along the diagonal and $d=\{40,80,120,160\}$ is considered. The off-diagonal blocks are zero, ensuring no inter-block correlation.
Results.
In critical high-discrepancy regimes (MI $\geq 20$ nats), ISA-DRE demonstrates unprecedented resilience: while DRE-$\infty$ collapses (MSE =283.52=283.52 at MI =20=20, NFE =2=2) and D3RE fails catastrophically at (MSE =500.04=500.04 at MI =40=40), ISA-DRE maintains sub-$1.0$ MSE ($0.72$ at MI =40=40, NFE =2=2). Crucially, with merely 2 function evaluations, our method achieves near-perfect estimation at MI =30=30 (MSE =0.18=0.18 at MI =30=30) where baselines require more than 5 times more computations to achieve comparable accuracy. This dimensional robustness is particularly evident at MI =40=40, where ISA-DRE sustains 3-orders-of-magnitude lower MSE than baselines ($0.72$ vs. $1215.69$ for DRE-$\infty$ at NFE =2=2), proving its capability to navigate high-discrepancy landscapes where traditional methods encounter pathological failures. These results validate that ISA-DRE’s novel secant-based transport planning successfully avoids density chasms by adaptively focusing computation on critical intermediate distributions.
C.4 Ablation Study
Ablation: Time Sampler.
The choice of time sampler critically influences performance in high-discrepancy conditions. VI sampling consistently delivers superior results, particularly at low NFEs. At MI =40=40 ($d=160$) with only $\text{NFE}=2$, VI sampling achieves $250\times$ lower MSE than Uniform sampling ($0.72$ vs $181.41$) and $30\times$ lower than LN sampling ($0.72$ vs $22.16$). This advantage stems from VI’s adaptive concentration of evaluations in high-discrepancy temporal regions, which becomes increasingly crucial as MI increases. While LN sampling shows moderate effectiveness at higher NFEs (e.g., MSE =6.78=6.78 at $d=160$, $\text{NFE}=5$ with LN/CIA), it remains substantially outperformed by VI sampling in extreme conditions (MSE =0.30=0.30 for VI/CIA vs $6.78$ for LN/CIA at same setting). Uniform sampling proves particularly vulnerable to density chasms, exhibiting catastrophic failures at MI =40=40 across multiple configurations (e.g., MSE =604.93=604.93 at $d=160$, $\text{NFE}=50$). These findings confirm adaptive time sampling is essential for navigating high-discrepancy landscapes.
(a) Asinh Mapping
(b) Additive Noise
(c) Half-Cube Map
(d) Edge-singular Gauss
Figure 7: Ablation studies on time samplers and secant-tangent supervision for MI estimations. Experiments are done on four geometrically pathological distributions. MSE and estimated MI are reported.
(a) Asinh Mapping
(b) Additive Noise
(c) Half-Cube Map
(d) Edge-singular Gauss
Figure 8: Ablation studies on time samplers for MI estimations. Experiments are done on four geometrically pathological distributions. MSE and estimated MI are reported.
Ablation: Secant-Tangent Supervision (STS).
STS configuration significantly impacts optimization stability, with secant-only ($0\%$ STS) demonstrating exceptional robustness in high-discrepancy regimes. At MI =40=40 ($d=160$) with $\text{NFE}=2$, $0\%$ STS achieves $77\times$ lower MSE than $100\%$ STS ($0.72$ vs $55.97$ for VI sampling). The CIA method provides an effective balance, particularly at higher NFEs: at $d=160$ with $\text{NFE}=50$, CIA with VI attains near-perfect estimation (MSE =0.01=0.01) while outperforming fixed STS ratios. Tangent-only ($100\%$ STS) consistently induces instability, causing catastrophic failures in multiple high-discrepancy settings (e.g., MSE =2804.96=2804.96 at $d=120$, $\text{NFE}=2$). Notably, CIA demonstrates dimensional adaptability: at $d=120$ (MI =30=30), it reduces MSE by $99.8\%$ compared to $100\%$ STS under VI sampling ($3.21$ vs $2804.96$ at $\text{NFE}=2$). These results establish that the CIA is paramount for stable high-discrepancy estimation, with endpoint-focused and CIA configurations proving most effective against density chasms.
(a) Asinh Mapping
(b) Additive Noise
(c) Half-Cube Map
(d) Edge-singular Gauss
Figure 9: Ablation studies on secant-tangent supervision for MI estimations. Experiments are done on four geometrically pathological distributions. MSE and estimated MI are reported.
NFE Method TS STS POWER GAS HEPMASS MINIBOONE BSDS300
2 DRE-$\infty$ VI $100\%$ $0.05\pm 1.84$ $4.37\pm 1.44$ $19.30\pm 1.31$ $41.55\pm 2.07$ $130.68\pm 4.17$
2 D3RE VI $100\%$ $3.57\pm 1.84$ $5.74\pm 15.28$ $23.90\pm 0.36$ $55.83\pm 9.36$ $149.53\pm 9.06$
2 ISA-DRE (ours) VI $0\%$ $3.26\pm 2.38$ $14.60\pm 3.50$ $12.35\pm 2.34$ $22.05\pm 3.19$ $\mathbf{-234.21}\pm\mathbf{15.30}$
2 ISA-DRE (ours) VI $50\%$ $3.66\pm 1.89$ $\mathbf{-14.95}\pm\mathbf{3.13}$ $\mathbf{12.07}\pm\mathbf{1.54}$ $3.50\pm 2.42 22$$20.94\pm 7.86$
2 ISA-DRE (ours) VI $100\%$ $3.87\pm 2.08$ $8.79\pm 0.82$ $13.28\pm 2.56$ $29.02\pm 7.47$ $197.74\pm 10.84$
2 ISA-DRE (ours) VI CIA $\mathbf{-5.07}\pm\mathbf{0.91}$ $10.63\pm 0.80$ $14.87\pm 3.22$ $\mathbf{20.05}\pm\mathbf{3.30}$ $196.15\pm 5.82$
2 ISA-DRE (ours) Uni. $0\%$ $0.54\pm 0.39$ $9.38\pm 5.72$ $\mathbf{14.55}\pm\mathbf{3.97}$ $20.78\pm 1.81$ $\mathbf{-89.48}\pm\mathbf{156.90}$
2 ISA-DRE (ours) Uni. $50\%$ $\mathbf{-1.82}\pm\mathbf{1.74}$ $\mathbf{-10.54}\pm\mathbf{2.97}$ $14.62\pm 3.33$ $19.37\pm 0.74$ $27.33\pm 235.53$
2 ISA-DRE (ours) Uni. $100\%$ $14.80\pm 0.53$ $5.66\pm 0.72$ $14.99\pm 3.13$ $22.77\pm 2.66$ $65.05\pm 41.39$
2 ISA-DRE (ours) Uni. CIA $0.11\pm 1.19$ $10.08\pm 2.91$ $16.18\pm 6.50$ $\mathbf{18.73}\pm\mathbf{0.70}$ $80.70\pm 12.44$
2 ISA-DRE (ours) LN $0\%$ $3.14\pm 0.24$ $\mathbf{-14.33}\pm\mathbf{0.98}$ $\mathbf{12.06}\pm\mathbf{0.47}$ $12.28\pm 2.85$ $\mathbf{-211.86}\pm\mathbf{16.11}$
[…]
2 ISA-DRE (ours) LN $100\%$ $\mathbf{-3.33}\pm\mathbf{0.23}$ $4.95\pm 6.43$ $13.53\pm 0.42$ $7.66\pm 0.67 11$$15.19\pm 22.68$
2 ISA-DRE (ours) LN CIA $1.61\pm 0.77$ $13.17\pm 1.15$ $13.40\pm 0.36$ $14.13\pm 1.07$ $205.23\pm 5.83$
5 DRE-$\infty$ VI $100\%$ $0.35\pm 0.50$ $3.63\pm 0.78$ $20.24\pm 0.47$ $20.90\pm 0.84$ $83.70\pm 1.35$
5 D3RE VI $100\%$ $26\pm 0.38 1.1$$15\pm 4.20$ $21.05\pm 0.52$ $43.11\pm 26.20$ $101.97\pm 1.67$
5 ISA-DRE (ours) VI $0\%$ $2.40\pm 1.24$ $10.27\pm 0.27$ $\mathbf{15.87}\pm\mathbf{1.06}$ $19.30\pm 2.41$ $\mathbf{-164.70}\pm\mathbf{10.48}$
5 ISA-DRE (ours) VI $50\%$ $2.30\pm 1.24$ $10.04\pm 0.85$ $15.91\pm 0.75$ $19.92\pm 1.34$ $153.30\pm 5.91$
5 ISA-DRE (ours) VI $00\% 1.1$$12\pm 0.44$ $2.42\pm 3.97$ $41.06\pm 35.22$ $29.08\pm 14.66$ $94.04\pm 2.23$
5 ISA-DRE (ours) VI CIA $\mathbf{-2.45}\pm\mathbf{0.39}$ $\mathbf{-10.56}\pm\mathbf{0.22}$ $16.60\pm 0.39$ $\mathbf{17.48}\pm\mathbf{0.38}$ $156.35\pm 7.45$
5 ISA-DRE (ours) Uni. $0\%$ $1.74\pm 0.26$ $10.26\pm 1.96$ $\mathbf{15.79}\pm\mathbf{1.06}$ $13.54\pm 0.47$ $152.30\pm 33.98$
5 ISA-DRE (ours) Uni. $50\%$ $\mathbf{-2.04}\pm\mathbf{0.78}$ $10.07\pm 1.44$ $15.80\pm 1.02$ $13.79\pm 0.04$ $139.38\pm 50.01$
5 ISA-DRE (ours) Uni. $100\%$ $1.78\pm 1.13$ $\mathbf{-10.83}\pm\mathbf{1.16}$ $17.41\pm 0.04$ $21.70\pm 5.46$ $\mathbf{-154.72}\pm\mathbf{2.25}$
5 ISA-DRE (ours) Uni. CIA $1.71\pm 0.28$ $10.04\pm 0.88$ $16.14\pm 1.66$ $\mathbf{13.04}\pm\mathbf{1.05}$ $138.61\pm 3.61$
5 ISA-DRE (ours) LN $0\%$ $\mathbf{-2.33}\pm\mathbf{0.00}$ $\mathbf{-9.97}\pm\mathbf{0.79}$ $16.34\pm 0.36$ $13.36\pm 1.44$ $153.54\pm 4.52$
5 ISA-DRE (ours) LN $50\%$ $1.74\pm 0.08$ $9.83\pm 0.24$ $16.76\pm 0.01$ $\mathbf{13.27}\pm\mathbf{1.16}$ $\mathbf{-153.76}\pm\mathbf{1.55}$
5 ISA-DRE (ours) LN $100\%$ $1.31\pm 0.43$ $9.41\pm 21.17 22$$2.47\pm 4.22$ $15.93\pm 0.41$ $53.08\pm 6.30$
5 ISA-DRE (ours) LN CIA $1.98\pm 0.04$ $9.58\pm 2.26$ $\mathbf{16.34}\pm\mathbf{1.01}$ $15.31\pm 1.63$ $150.16\pm 2.03$
10 DRE-$\infty$ VI $100\%$ $0.03\pm 0.17$ $4.34\pm 0.60$ $20.43\pm 0.52$ $20.57\pm 0.93$ $87.65\pm 2.24$
10 D3RE VI $100\%$ $0.49\pm 0.39$ $3.27\pm 2.00$ $20.30\pm 0.55$ $42.65\pm 26.87$ $102.01\pm 2.43$
10 ISA-DRE (ours) VI $0\%$ $0.94\pm 1.26$ $7.59\pm 0.54$ $\mathbf{17.92}\pm\mathbf{1.15}$ $18.45\pm 1.92$ $139.62\pm 7.72$
10 ISA-DRE (ours) VI $50\%$ $0.70\pm 1.24$ $7.59\pm 0.83$ $17.93\pm 0.84$ $19.11\pm 1.39$ $128.78\pm 2.91$
10 ISA-DRE (ours) VI $100\%$ $\mathbf{-1.97}\pm\mathbf{0.40}$ $3.98\pm 8.42$ $18.89\pm 2.61$ $29.40\pm 16.32$ $55.54\pm 8.25$
10 ISA-DRE (ours) VI CIA $1.17\pm 0.08$ $\mathbf{-9.66}\pm\mathbf{0.01}$ $18.23\pm 0.36$ $\mathbf{17.29}\pm\mathbf{0.48}$ $\mathbf{-162.14}\pm\mathbf{3.98}$
10 ISA-DRE (ours) Uni. $0\%$ $0.69\pm 0.30$ $8.53\pm 1.25$ $17.66\pm 0.60$ $13.19\pm 0.47$ $\mathbf{-160.03}\pm\mathbf{10.48}$
10 ISA-DRE (ours) Uni. $50\%$ $\mathbf{-0.92}\pm\mathbf{0.59}$ $8.15\pm 1.11$ $\mathbf{17.62}\pm\mathbf{0.70}$ $13.73\pm 0.13$ $155.48\pm 16.12$
10 ISA-DRE (ours) Uni. $100\%$ $0.52\pm 0.39$ $\mathbf{-8.82}\pm\mathbf{1.32}$ $18.41\pm 0.01$ $51.34\pm 51.17$ $110.97\pm 33.15$
10 ISA-DRE (ours) Uni. CIA $0.80\pm 0.15$ $7.34\pm 0.14$ $17.70\pm 0.77$ $\mathbf{13.04}\pm\mathbf{0.96}$ $154.75\pm 2.27$
10 ISA-DRE (ours) LN $0\%$ $\mathbf{-0.92}\pm\mathbf{0.19}$ $6.28\pm 0.91$ $\mathbf{18.46}\pm\mathbf{0.00}$ $13.47\pm 1.47$ $131.71\pm 2.71$
10 ISA-DRE (ours) LN $50\%$ $0.58\pm 0.47$ $6.51\pm 0.19$ $18.86\pm 0.63$ $13.32\pm 0.84$ $129.67\pm 1.89$
10 ISA-DRE (ours) LN $100\%$ $0.38\pm 0.21$ $57.55\pm 25.63$ $38.45\pm 24.31$ $22.08\pm 4.13$ $29.70\pm 14.58$
10 ISA-DRE (ours) LN CIA $0.65\pm 0.45$ $\mathbf{-8.53}\pm\mathbf{0.21}$ $18.48\pm 0.98$ $\mathbf{12.98}\pm\mathbf{0.45}$ $\mathbf{-148.53}\pm\mathbf{3.16}$
50 DRE-$\infty$ VI $100\%$ $0.25\pm 0.28$ $4.33\pm 0.71$ $20.67\pm 0.57$ $20.97\pm 0.51$ $90.24\pm 2.14$
50 D3RE VI $100\%$ $0.89\pm 0.33$ $3.16\pm 0.62$ $20.05\pm 0.35$ $42.73\pm 26.78$ $78.26\pm 0.96$
50 ISA-DRE (ours) VI $0\%$ $0.21\pm 0.97$ $5.82\pm 0.37$ $19.27\pm 0.95$ $8.36\pm 1.32 11$$19.65\pm 5.88$
50 ISA-DRE (ours) VI $50\%$ $0.03\pm 0.99$ $5.80\pm 0.61$ $19.34\pm 0.72$ $8.57\pm 1.33 11$$10.75\pm 1.97$
50 ISA-DRE (ours) VI $100\%$ $\mathbf{-2.14}\pm\mathbf{0.66}$ $4.42\pm 10.17$ $\mathbf{19.25}\pm\mathbf{0.31}$ $31.62\pm 20.22$ $50.47\pm 19.94$
50 ISA-DRE (ours) VI CIA $0.61\pm 0.37$ $\mathbf{-8.19}\pm\mathbf{0.12}$ $19.46\pm 0.41$ $\mathbf{17.34}\pm\mathbf{0.41}$ $\mathbf{-150.54}\pm\mathbf{1.71}$
50 ISA-DRE (ours) Uni. $\% 0.0$$03\pm 0.33$ $6.39\pm 0.95$ $19.07\pm 0.37$ $13.44\pm 0.40$ $145.34\pm 2.54$
50 ISA-DRE (ours) Uni. $50\%$ $0.10\pm 0.47$ $5.92\pm 1.01$ $\mathbf{19.00}\pm\mathbf{0.49}$ $14.13\pm 0.15$ $142.86\pm 3.50$
50 ISA-DRE (ours) Uni. $100\%$ $\mathbf{-0.17}\pm\mathbf{0.35}$ $\mathbf{-6.49}\pm\mathbf{0.98}$ $19.03\pm 0.81$ $61.43\pm 108.42$ $85.03\pm 36.49$
50 ISA-DRE (ours) Uni. CIA $0.14\pm 0.10$ $5.50\pm 0.59$ $19.08\pm 0.49$ $\mathbf{13.41}\pm\mathbf{0.70}$ $\mathbf{-147.21}\pm\mathbf{0.54}$
50 ISA-DRE (ours) LN $0\%$ $0.37\pm 0.11$ $5.12\pm 0.70$ $\mathbf{19.47}\pm\mathbf{0.03}$ $2.95\pm 1.02 11$$12.82\pm 3.25$
50 ISA-DRE (ours) LN $50\%$ $0.07\pm 0.14$ $5.39\pm 0.26$ $19.81\pm 0.49$ $2.75\pm 0.60 11$$11.92\pm 1.62$
50 ISA-DRE (ours) LN $100\%$ $\mathbf{-0.80}\pm\mathbf{0.27}$ $52.08\pm 26.42$ $42.23\pm 31.25$ $34.48\pm 14.49$ $31.10\pm 28.62$
50 ISA-DRE (ours) LN CIA $0.03\pm 0.30$ $\mathbf{-6.24}\pm\mathbf{0.46}$ $19.66\pm 0.84$ $\mathbf{12.58}\pm\mathbf{0.64}$ $\mathbf{-133.41}\pm\mathbf{2.05}$
Table 4: Density estimation results on five real-world tabular datasets with complex, non-Gaussian structures. Values indicate negative log-likelihood (NLL; lower is better), reported as $\text{mean}\pm\text{std}$ over 3 runs. Results are shown across varying function evaluations ($\text{NFE}\in\{2,5,10,50\}$), time samplers (TS; see Sec. 4.3), and secant-tangent supervisions (STS; see Sec. 4.3). Bold entries mark the best mean NLL for each NFE-TS-dataset setting.
$d=40$ (MI =10=10) $d=80$ (MI =20=20) $d=120$ (MI =30=30) $d=160$ (MI =40=40)
NFE Method TS STS Est. MI MSE Est. MI MSE Est. MI MSE Est. MI MSE
22 DRE-$\infty$ VI $100\%$ $1.40\pm 0.01$ $73.91$ $3.16\pm 0.01$ $283.52$ $5.21\pm 0.01$ $614.62$ $5.13\pm 0.02$ $1215.69$
22 D3RE VI $00\% 11$$1.61\pm 0.08$ $2.58$ $21.91\pm 0.08$ $3.65$ $27.51\pm 0.07$ $6.21$ $17.64\pm 0.17$ $500.04$
22 ISA-DRE (ours) VI $0\%$ $10.13\pm 0.11$ $0.03$ $19.59\pm 0.07$ $\mathbf{0.18}$ $27.99\pm 0.05$ $4.03$ $40.85\pm 0.06$ $\mathbf{0.72}$
22 ISA-DRE (ours) VI $50\%$ $0\pm 0.11 0.01\mathbf{0.01}$ $18.64\pm 0.07$ $1.84$ $27.18\pm 0.09$ $7.99$ $43.10\pm 0.06$ $9.62$
22 ISA-DRE (ours) VI $100\%$ $7.97\pm 0.03$ $4.14$ $26.08\pm 0.04$ $36.97$ $82.96\pm 0.02$ $2804.96$ $32.52\pm 0.10$ $55.97$
22 ISA-DRE (ours) VI CIA $10.68\pm 0.08$ $0.47$ $21.19\pm 0.08$ $1.42$ $31.79\pm 0.08$ $\mathbf{3.21}$ $41.33\pm 0.10$ $1.78$
22 ISA-DRE (ours) Uni. $0\%$ $7.92\pm 0.13$ $4.33$ $15.20\pm 0.12$ $23.08$ $21.23\pm 0.14$ $77.01$ $26.53\pm 0.04$ $181.41$
22 ISA-DRE (ours) Uni. $50\%$ $7.67\pm 0.13$ $5.44$ $17.46\pm 0.13$ $\mathbf{6.45}$ $9.26\pm 0.19 11$5.29115.29 $28.32\pm 0.07$ $136.50$
22 ISA-DRE (ours) Uni. $00\% 11$$1.51\pm 0.12$ $2.28$ $28.72\pm 0.12$ $76.12$ $27.95\pm 0.13$ $4.22$ $28.76\pm 0.05$ $126.23$
22 ISA-DRE (ours) Uni. CIA $8.87\pm 0.13$ $\mathbf{1.31}$ $15.80\pm 0.10$ $17.63$ $28.97\pm 0.16$ $\mathbf{1.09}$ $29.27\pm 0.07$ $\mathbf{115.04}$
22 ISA-DRE (ours) LN $0\%$ $11.71\pm 0.07$ $\mathbf{2.93}$ $24.56\pm 0.04$ $20.78$ $32.94\pm 0.04$ $8.65$ $44.71\pm 0.02$ $22.16$
22 ISA-DRE (ours) LN $50\%$ $11.83\pm 0.06$ $3.35$ $23.88\pm 0.05$ $15.04$ $33.39\pm 0.06$ $11.47$ $43.75\pm 0.02$ $14.08$
22 ISA-DRE (ours) LN $100\%$ $18.08\pm 0.07$ $65.31$ $24.61\pm 0.05$ $21.28$ $34.74\pm 0.04$ $22.47$ $43.23\pm 0.01$ $10.43$
22 ISA-DRE (ours) LN CIA $11.86\pm 0.05$ $3.45$ $21.61\pm 0.04$ $\mathbf{2.60}$ $31.74\pm 0.04$ $\mathbf{3.02}$ $40.91\pm 0.03$ $\mathbf{0.83}$
55 DRE-$\infty$ VI $100\%$ $8.31\pm 0.05$ $2.86$ $17.34\pm 0.04$ $7.09$ $24.97\pm 0.05$ $25.29$ $31.61\pm 0.06$ $70.36$
55 D3RE VI $100\%$ $9.91\pm 0.04$ $\mathbf{0.01}$ $19.46\pm 0.04$ $\mathbf{0.29}$ $27.26\pm 0.03$ $7.50$ $31.24\pm 0.05$ $76.80$
55 ISA-DRE (ours) VI $0\%$ $11.29\pm 0.05$ $1.67$ $22.59\pm 0.04$ $6.70$ $32.78\pm 0.03$ $7.74$ $44.85\pm 0.03$ $23.52$
55 ISA-DRE (ours) VI $50\%$ $11.49\pm 0.05$ $2.23$ $22.59\pm 0.04$ $6.69$ $33.51\pm 0.05$ $12.30$ $45.57\pm 0.03$ $31.02$
55 ISA-DRE (ours) VI $100\%$ $5.77\pm 0.03$ $17.88$ $18.75\pm 0.03$ $1.57$ $67.07\pm 0.03$ $1374.40$ $41.40\pm 0.04$ $1.96$
55 ISA-DRE (ours) VI CIA $11.92\pm 0.06$ $3.68$ $22.47\pm 0.04$ $6.12$ $32.13\pm 0.04$ $\mathbf{4.56}$ $39.45\pm 0.05$ $\mathbf{0.30}$
55 ISA-DRE (ours) Uni. $0\%$ $1.06\pm 0.06 1.1$21.12 $22.23\pm 0.06$ $4.99$ $31.27\pm 0.05$ $1.61$ $41.83\pm 0.02$ $3.36$
55 ISA-DRE (ours) Uni. $50\%$ $10.82\pm 0.06$ $0.68$ $23.07\pm 0.06$ $9.45$ $31.20\pm 0.06$ $\mathbf{1.45}$ $41.86\pm 0.02$ $3.46$
55 ISA-DRE (ours) Uni. $100\%$ $10.37\pm 0.05$ $\mathbf{0.14}$ $30.53\pm 0.05$ $110.91$ $41.92\pm 0.03$ $142.07$ $41.41\pm 0.12$ $2.00$
55 ISA-DRE (ours) Uni. CIA $11.74\pm 0.07$ $3.03$ $21.99\pm 0.05$ $\mathbf{3.98}$ $31.22\pm 0.06$ $1.50$ $40.34\pm 0.04$ $\mathbf{0.12}$
55 ISA-DRE (ours) LN $0\%$ $12.05\pm 0.05$ $4.21$ $22.99\pm 0.03$ $8.94$ $34.98\pm 0.03$ $24.81$ $44.06\pm 0.03$ $16.50$
55 ISA-DRE (ours) LN $50\%$ $12.18\pm 0.05$ $4.77$ $22.68\pm 0.03$ $7.16$ $34.38\pm 0.02$ $19.21$ $42.60\pm 0.03$ $\mathbf{6.78}$
55 ISA-DRE (ours) LN $100\%$ $13.68\pm 0.04$ $13.52$ $29.54\pm 0.02$ $90.93$ $67.95\pm 0.01$ $1439.98$ $60.91\pm 0.01$ $437.32$
55 ISA-DRE (ours) LN CIA $11.74\pm 0.04$ $\mathbf{3.01}$ $21.53\pm 0.03$ $\mathbf{2.33}$ $31.27\pm 0.04$ $\mathbf{1.63}$ $42.82\pm 0.05$ $7.98$
$10$ DRE-$\infty$ VI $100\%$ $9.48\pm 0.06$ $0.27$ $19.27\pm 0.04$ $0.54$ $28.37\pm 0.05$ $2.66$ $37.34\pm 0.06$ $7.08$
$10$ D3RE VI $100\%$ $10.13\pm 0.04$ $\mathbf{0.02}$ $20.45\pm 0.03$ $\mathbf{0.21}$ $27.22\pm 0.03$ $7.72$ $32.27\pm 0.04$ $59.70$
$10$ ISA-DRE (ours) VI $0\%$ $10.45\pm 0.04$ $0.20$ $21.14\pm 0.04$ $1.29$ $30.86\pm 0.03$ $0.75$ $41.64\pm 0.03$ $2.68$
$10$ ISA-DRE (ours) VI $50\%$ $10.70\pm 0.04$ $0.49$ $21.39\pm 0.03$ $1.94$ $32.00\pm 0.04$ $3.99$ $42.19\pm 0.03$ $4.79$
$10$ ISA-DRE (ours) VI $100\%$ $5.90\pm 0.03$ $16.82$ $15.26\pm 0.03$ $22.45$ $73.87\pm 0.03$ $1924.57$ $37.11\pm 0.04$ $8.33$
$10$ ISA-DRE (ours) VI CIA $1.09\pm 0.05 1.1$91.19 $20.92\pm 0.03$ $0.86$ $30.22\pm 0.04$ $\mathbf{0.05}$ $40.74\pm 0.04$ $\mathbf{0.54}$
$10$ ISA-DRE (ours) Uni. $0\%$ $10.68\pm 0.06$ $0.47$ $21.76\pm 0.05$ $3.09$ $30.49\pm 0.04$ $0.25$ $41.00\pm 0.02$ $0.99$
$10$ ISA-DRE (ours) Uni. $50\%$ $10.61\pm 0.05$ $\mathbf{0.37}$ $22.26\pm 0.05$ $5.12$ $30.28\pm 0.04$ $0.08$ $40.33\pm 0.02$ $\mathbf{0.11}$
$10$ ISA-DRE (ours) Uni. $100\%$ $12.30\pm 0.06$ $5.31$ $31.83\pm 0.05$ $139.89$ $102.81\pm 0.04$ $5300.98$ $46.62\pm 0.10$ $43.86$
$10$ ISA-DRE (ours) Uni. CIA $11.40\pm 0.06$ $1.96$ $21.46\pm 0.04$ $\mathbf{2.12}$ $30.24\pm 0.05$ $\mathbf{0.06}$ $39.48\pm 0.03$ $0.27$
$10$ ISA-DRE (ours) LN $0\%$ $11.49\pm 0.05$ $2.23$ $22.09\pm 0.03$ $4.38$ $32.56\pm 0.03$ $6.57$ $39.84\pm 0.03$ $0.03$
$10$ ISA-DRE (ours) LN $50\%$ $11.62\pm 0.05$ $2.61$ $21.97\pm 0.03$ $3.88$ $31.33\pm 0.02$ $1.77$ $38.68\pm 0.03$ $1.76$
$10$ ISA-DRE (ours) LN $100\%$ $26.91\pm 0.04$ $286.02$ $55.68\pm 0.04$ $1273.42$ $39.60\pm 0.02$ $92.13$ $32.71\pm 0.03$ $53.08$
$10$ ISA-DRE (ours) LN CIA $10.84\pm 0.05$ $\mathbf{0.71}$ $20.33\pm 0.04$ $\mathbf{0.11}$ $30.71\pm 0.05$ $\mathbf{0.51}$ $39.89\pm 0.07$ $\mathbf{0.02}$
$50$ DRE-$\infty$ VI $100\%$ $9.84\pm 0.06$ $0.03$ $19.81\pm 0.04$ $0.04$ $29.31\pm 0.06$ $0.48$ $38.06\pm 0.07$ $3.77$
$50$ D3RE VI $100\%$ $7\pm 0.04 0.010.01$ $20.30\pm 0.03$ $0.09$ $27.01\pm 0.03$ $8.94$ $32.37\pm 0.04$ $58.19$
$50$ ISA-DRE (ours) VI $0\%$ $9.96\pm 0.04$ $\mathbf{0.00}$ $20.22\pm 0.03$ $0.05$ $29.98\pm 0.02$ $\mathbf{0.00}$ $40.49\pm 0.03$ $0.25$
$50$ ISA-DRE (ours) VI $50\%$ $10.26\pm 0.04$ $0.07$ $20.58\pm 0.03$ $0.34$ $31.26\pm 0.04$ $1.58$ $40.88\pm 0.03$ $0.77$
$50$ ISA-DRE (ours) VI $100\%$ $9.83\pm 0.03$ $0.03$ $20.60\pm 0.03$ $0.36$ $28.89\pm 0.04$ $1.23$ $40.74\pm 0.04$ $0.54$
$50$ ISA-DRE (ours) VI CIA $10.90\pm 0.05$ $0.81$ $20.01\pm 0.02$ $\mathbf{0.00}$ $30.05\pm 0.01$ $0.00$ $39.91\pm 0.04$ $\mathbf{0.01}$
$50$ ISA-DRE (ours) Uni. $0\%$ $10.36\pm 0.05$ $0.14$ $21.28\pm 0.05$ $1.65$ $29.72\pm 0.03$ $0.08$ $39.83\pm 0.02$ $\mathbf{0.03}$
$50$ ISA-DRE (ours) Uni. $50\%$ $10.31\pm 0.05$ $\mathbf{0.10}$ $21.44\pm 0.05$ $2.07$ $29.45\pm 0.03$ $0.31$ $38.97\pm 0.02$ $1.07$
$50$ ISA-DRE (ours) Uni. $100\%$ $10.68\pm 0.06$ $0.46$ $48.40\pm 0.04$ $806.36$ $60.76\pm 0.05$ $946.06$ $15.41\pm 0.14$ $604.93$
$50$ ISA-DRE (ours) Uni. CIA $11.14\pm 0.06$ $1.31$ $21.13\pm 0.04$ $\mathbf{1.29}$ $30.17\pm 0.04$ $\mathbf{0.03}$ $39.32\pm 0.03$ $0.46$
$50$ ISA-DRE (ours) LN $0\%$ $11.23\pm 0.05$ $1.52$ $21.27\pm 0.03$ $1.61$ $32.47\pm 0.03$ $6.08$ $38.56\pm 0.03$ $2.07$
$50$ ISA-DRE (ours) LN $50\%$ $11.27\pm 0.05$ $1.63$ $21.24\pm 0.03$ $1.53$ $30.56\pm 0.02$ $0.31$ $37.18\pm 0.04$ $7.97$
$50$ ISA-DRE (ours) LN $100\%$ $17.79\pm 0.07$ $60.61$ $32.22\pm 0.07$ $149.25$ $36.75\pm 0.03$ $45.62$ $35.87\pm 0.03$ $17.02$
$50$ ISA-DRE (ours) LN CIA $10.57\pm 0.05$ $\mathbf{0.32}$ $20.35\pm 0.04$ $\mathbf{0.12}$ $30.41\pm 0.06$ $\mathbf{0.17}$ $39.69\pm 0.08$ $\mathbf{0.11}$
Table 5: Mutual information estimation results under high-discrepancy (MI $\in\{10,20,30,40\}$ nats). We report estimated MI (mean $\pm$ std over 5 seeds) and MSE for different function evaluations (NFE $\in\{2,5,10,50\}$), time samplers (TS, see Sec. 4.3), and secant-tangent supervision (STS, see Sec. 4.3). Bold MSE indicates best performance for each NFE-TS-$d$ combination. ISA-DRE consistently succeeds where others collapse.